### v1

# InSyn: Modeling Complex Interactions for Pedestrian Trajectory Prediction

Abstract
Accurate pedestrian trajectory prediction is crucial for intelligent applications, yet it remains highly challenging due to the complexity of interactions among pedestrians. Previous methods have primarily relied on relative positions to model pedestrian interactions; however, they tend to overlook specific interaction patterns such as paired walking or conflicting behaviors, limiting the prediction accuracy in crowded scenarios. To address this issue, we propose InSyn (Interaction-Synchronization Network), a novel Transformer-based model that explicitly captures diverse interaction patterns (e.g., walking in sync or conflicting) while effectively modeling direction-sensitive social behaviors. Additionally, we introduce a training strategy termed Seq-Start of Seq (SSOS), designed to alleviate the common issue of initial-step divergence in numerical time-series prediction. Experiments on the ETH and UCY datasets demonstrate that our model outperforms recent baselines significantly, especially in high-density scenarios. Furthermore, the SSOS strategy proves effective in improving sequential prediction performance, reducing the initial-step prediction error by approximately 6.58%.
[…]
7556
1 Introduction
Pedestrian trajectory prediction plays a crucial role in various fields such as autonomous driving [21], robotic navigation [4], and smart cities [5]. Accurately forecasting a pedestrian’s future trajectory not only helps intelligent systems better understand human behavior but also significantly enhances the safety and efficiency of human-machine interaction. However, beyond an individual’s own motion, interactions among pedestrians are often highly complex, making trajectory prediction a challenging task. In a given scene, pedestrians may respond to nearby individuals in different ways. For instance, the trajectories of two pedestrians may exhibit conflict. Alternatively, they may show weak or no influence on each other, such as when walking in sync or as part of a group. Notably, even in walk-in-sync scenarios, subtle interactions may still occur. Effectively modeling these intricate interaction patterns, especially in high-density environments, remains a major challenge in pedestrian trajectory prediction.
Figure 1: Comparison of Interaction Modeling: Previous Methods vs. Our Approach. F represents the interaction affect between pedestrians. In traditional approaches (top), all neighbors of the agent are treated as being in the same state, with interactions solely based on the distance between the agent and its neighbors. This simplistic assumption may lead to overestimating the influence of fellow pedestrians. Our method (bottom) introduces a more refined modeling strategy by considering the specific states of neighboring pedestrians (e.g., In Sync or Conflict), providing a more nuanced understanding of the neighboring interaction.
Recently, extensive studies have investigated pedestrian interactions [1, 32, 25, 43], using approaches like social pooling layers [1, 7], social force mechanism [43], and graph neural networks (GNNs) [41, 19]. However, a common characteristic of these methods is that they rely on relative positions between an agent and its neighbors to model influence, without considering the specific interaction relationships between them. This limitation restricts the capacity of previous models to capture the dynamics of real-world interactions, such as walking in sync or avoiding collisions, which are crucial for accurately predicting trajectory in crowded scenario. Prior approaches treat all neighboring pedestrians uniformly, failing to differentiate patterns such as In Sync and Conflict (see Figure 1). This oversimplification may leads to overfitting on sparse datasets and underperforming in complex social scenarios. For instance, in situations where pedestrians walk in sync (e.g., friends walking side-by-side), failing to account for this coordinated behavior may lead to overlapping trajectories during prediction. Similarly, in conflict scenarios (e.g., two pedestrians approaching each other head-on or navigating narrow passages), the lack of explicit modeling for competitive or avoidance behaviors often results in unrealistic trajectories that either overestimate collision risk or underestimate the necessity for trajectory alteration. These limitations become particularly pronounced in crowded or dynamic environments.
To overcome this issue, we need a method that explicitly identifies and captures these interaction patterns, enabling a deeper understanding of the diverse pedestrian interactions, as shown in Figure 1.
Although pedestrian trajectories are highly stochastic, they usually have a clear goal [28]. By focusing on the target goal, the performance of the model can be significantly enhanced. Experimental results from Giuliari et al. [9] demonstrate that the attention mechanism is well-suited for processing sequential data with irregular time intervals (e.g., observed trajectories and goal), effectively capturing long-term dependencies in pedestrian motion and providing reliable goal-driven functionality.
This paper proposes the InSyn, a Transformer-based model. The model consists of three components: (1) Interaction Encoder: designed to explicitly extract interaction information and integrates interaction features, goal-driven behavior, and observed trajectories through the self-attention mechanism. (2) Trajectory Generator: forms an attention mechanism in conjunction with the Interaction Encoder, and incorporates the proposed SSOS strategy to alleviate divergence at the initial prediction step. (3) Seq-CVAE Goal Sampler: a conditional generative model specifically designed for sequential prediction, used for goals sampling.
We evaluate our model on the well-established pedestrian trajectory prediction datasets ETH [22] and UCY [16]. The experimental results demonstrate that our model outperforms recent methods in key metrics, particularly in scenarios with complex interactions. We further conduct ablation studies to show the superiority of interaction modeling and SSOS strategy.
Figure 2: Left is the illustration of the input $S_{0:\tau}$. At each time step, the walking state $S_{k}$ comprises the 2D coordinates $(x,y)$ and interaction information $N_{k}$. Right demonstrates the scenarios of In Sync, Conflict and No Neighbor state.
In summary, our work makes two key contributions:
  • We propose a pedestrian interaction modeling approach that explicitly identifies and leverages specific interaction patterns, achieving a significant improvement in average ADE compared to the previous baselines and offering more comprehensive insights into socially-aware motion patterns.
  • For numerical time-series prediction tasks, we introduce a novel training strategy for the Transformer encoder-decoder architecture. This strategy, termed SSOS, mitigates the divergence in the first prediction step, thereby reducing error accumulation and improving overall performance.
2 Related Work
2.1 Pedestrian Trajectory Prediction
Pedestrian trajectory prediction methods can be categorized into traditional machine learning approaches and deep learning-based approaches. Traditional methods, such as Markov decision processes [30] and Gaussian distributions [35], are efficient in handling simple scenarios. However, they often struggle with complex data distributions and exhibit limited generalization capabilities. In contrast, deep learning approaches have shown significant improvements in modeling complex patterns [13, 8, 6, 36]. Long Short-Term Memory (LSTM) network [11], a classic time-series model, is applied in this field. Social-LSTM [1] introduced a pooling mechanism to incorporate interactions into trajectory prediction. Yang et al. [40] proposed a variant LSTM structure for information sharing within scenes, while SR-LSTM [44] combined neighbor intentions and message passing to enhance prediction accuracy. These work leverage the strong temporal feature extraction capabilities of LSTM.
In recent years, Transformer have gained popularity in trajectory prediction due to their ability to capture long-range dependencies. Liu et al. [20] combined CNN and Transformer, using attention mechanism to process interaction information. MRGTraj [23] integrated a Transformer encoder with a novel decoder, achieving trajectory prediction through a mapping-refining-generating structure. CITraNet [3] significantly improved prediction accuracy by introducing an innovative Transformer-based Gumbel distribution network.
Building on these advancements, our research combines the strengths of LSTM and Transformer. By leveraging the temporal feature extraction capabilities of LSTM, we enhance the Transformer’s attention mechanism with processed temporal interactions features, thereby improving prediction performance.
2.2 Pedestrian Interaction Modeling
Pedestrian interaction is a key factor influencing prediction accuracy. Pedestrian motion is not only driven by individual goals but also affected by interactions with surrounding pedestrians.
Current mainstream methods for interaction modeling include social pooling layers [1, 37], GNNs [19, 41], and attention mechanisms [18, 38]. Among these, the social attention module aggregates interaction information by analyzing correlations between pedestrian motion and future trajectories [37]. Yang et al. [38] introduced a social graph attention mechanism combined with a pseudo-oracle predictor to capture social interactions and intention states, enhancing prediction accuracy. Transformer have been utilized in pedestrian interaction modeling [34, 39, 12, 2, 27]. For example, Yuan et al. [42] proposed AgentFormer that learns spatiotemporal interaction embeddings from sequential trajectory features. Similarly, Yang et al. [39] leveraged GNNs and Transformer to model spatial and temporal dependencies. To address future interaction modeling, Amirloo et al. [2] employed the self-attention mechanism of Transformer to model pedestrian interactions and considered future states autoregressively during decoding to avoid trajectory conflicts.
However, most of these methods fail to consider the specific interaction relationships among pedestrians, such as walking In Sync or engaging in Conflict. To address this limitation, we propose a model that explicitly captures specific types of interaction features.
3 Problem Formulation
In this study, pedestrian trajectory prediction is formulated as follows: given the walking states $S_{0:\tau}$ of a certain agent during the observation time steps $\mathord{0:\tau}$, the model outputs the predicted trajectory $\widehat{pos}_{\tau+1:T}$ of the agent for the future time steps $\mathord{\tau+1:T}$. The input walking states of agent $p$ are denoted as $\mathcal{S}^{p}=\left(S_{0}^{p},S_{1}^{p},\dots,S_{\tau}^{p}\right)$, where $S_{t}^{p}$ represents the walking state at time step $t\ (0\leq t\leq\tau)$. Each walking state $S_{t}^{p}$ consists of the agent’s position $pos_{t}=(x_{t},y_{t})$ in 2D space and the pedestrian interaction information $N_{t}$. Since interactions are directional, we adopt a region partition strategy around each agent. The interaction information $N_{t}$ is defined as
$N_{t}=\left[\left(f_{t}^{lu},d_{t}^{lu}\right),\left(f_{t}^{ru},d_{t}^{ru}\right),\left(f_{t}^{ld},d_{t}^{ld}\right),\left(f_{t}^{rd},d_{t}^{rd}\right)\right]$ (1)
where $f_{t}^{lu},f_{t}^{ru},f_{t}^{ld},f_{t}^{rd}$ represent the interaction states in the left-up, right-up, left-down, and right-down regions respectively. These patterns include No Neighbor, In Sync, and Conflict, as shown in Figure 2 .
Specifically, if the nearest neighbor in a given region at time $t$ is the same as the one at time $t-1$, the interaction state is classified as In Sync; otherwise, it is marked as Conflict. In cases of No Neighbor or In Sync, the interaction influence of the neighbor on the agent’s future trajectory may be negligible. However, if a new pedestrian suddenly enters in a region, referred as Conflict, it is more likely to impact the agent’s trajectory. The terms $d_{t}^{lu},d_{t}^{ru},d_{t}^{ld},d_{t}^{rd}$ indicate the distances to the nearest neighbor in each region. A region with no pedestrians is classified as No Neighbor, and a large distance value is assigned to indicate minimal influence. The goal of this study is to train a generative model
$m_{\theta}\left(\widehat{pos}_{\tau+1:T}\mid S_{0:\tau}\right)$ (2)
where $\theta$ represents the model parameters. The model aims to predict the distribution of pedestrian’s future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
Figure 3: Overview of the InSyn framework for trajectory prediction. Our model consists of three key modules: (1) Interaction Encoder, (2) Trajectory Generator, and (3) Seq-CVAE. The input observed walking state includes the agent’s trajectory positions $pos_{0:\tau}$ and its interaction information $N_{0:\tau}$ within the obserbed time $0:\tau$.
4 Methodology
4.1 Framework Design
In this section, we propose InSyn, a Transformer-based model designed to identify and model spefific social interaction. The model consists of three components: (1) Interaction Encoder: Extracts interaction features and enables goal-driven prediction. (2) Trajectory Generator: Employing SSOS strategy to mitigate initial prediction step divergence. (3) Seq-CVAE Goal Sampler: A generator for sequential data that utilizes high-dimensional latent variables for goal sampling. The framework of InSyn is illustrated in Figure 3, and the Seq-CVAE is depicted in Figure 5.
We first discuss how our model captures specific interactions and achieves goal-driven functionality. We then elaborate on the motivation behind the SSOS strategy and, finally, the design of the Seq-CVAE module.
4.2 Interaction Encoder
The Interaction Encoder integrates observed trajectory, interaction information, and sampled goal to extract spatiotemporal features. These features are utilized for goal-driven and interaction-sensitive trajectory prediction.
4.2.1 Goal-Driven
Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  4.3) is combined with the position information through concatenation. Notably, since self-attention cannot inherently capture positional information, positional encoding is used to encode the input positions. This encoding is based on the sequence positions of the inputs. Therefore, before concatenating $pos_{0:\tau}^{p}$ with the sampled goal $\widehat{pos}_{T}$, padding needs to be inserted between them to ensure that the goal is assigned the corresponding positional encoding.
4.2.2 Interaction
The complex interactions among pedestrians are extracted through a neighbor encoder module follwed by an LSTM. The agent $p$’s social interaction information, represented as $\mathcal{N}^{p}=\left(N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}\right)$, are encoded using the neighbor encoder. The design of the neighbor encoder includes three components: the embedding layer, the Distance Gate, and the Social Pooling Layer. The embedding layer maps three interaction states (No Neighbor, In Sync, Conflict) into corresponding representations. These representations are then combined with the processed distance $d$ through element-wise multiplication. The process of $d$ is through the Distance Gate: $\mathrm{Gate}(d)=\sigma(\mathbf{W}\cdot d+\mathbf{b})$, where $\sigma(z)=\frac{1}{1+e^{-z}}$ and $\mathbf{W},\mathbf{b}$ represent the parameters of linear layer. The Distance Gate is similar to the gate mechanism in LSTM [11]. It maps $d$ to a weight in the range [0,1]. This weight reflects distance-based intensity, enhancing the spatial characteristics of pedestrian interactions.
After forming four regional interaction representations, they are passed through the Social Pooling Layer, which applies a max-pooling function across the region dimension. This enables the model to focus on the most influential interactions while suppressing less relevant or redundant information. The complete architecture of the Neighbor Encoder is illustrated in Figure 3.
The social interaction information $N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}$ for agent $p$ is processed by the Neighbor Encoder to produce the output $E_{0}^{p},E_{1}^{p},...,E_{\tau}^{p}$, which is then fed into the LSTM. The hidden state of LSTM at the final observed time step $h_{\tau}$ contains all temporal information, thus no need for positional encoding. Consequently, $h_{\tau}$ is directly concatenated with the positionally encoded representation (see Figure 3).
4.3 SSOS: Seq-Start of Seq
The Trajectory Generator incorporates the SSOS Strategy into the Transformer Decoder [29] to alleviate initial prediction divergence. The Transformer’s decoder relies on the Start of Sequence (SOS) as the initial input when generating the first output. In Natural Language Processing (NLP), the Begin-of-Sentence token $bos\right>$ is commonly used as the SOS, which aligns with the semantic structure of text. However, in sequential trajectory prediction, the selection of SOS should consider the spatial information; otherwise, it may introduce noise. Giuliari et al. [9] set the SOS to $(0,0)$ for trajectories prediction. While $(0,0)$ represents a fixed position in spatial coordinates, using it as the SOS may introduce noise and mislead the model due to its lack of alignment with the observed trajectory. Therefore, it is reasonable to consider using the position at the last observed time step ,$\widehat{pos}_{\tau}$, as the SOS token of decoder.
Nevertheless, we note that when the Transformer’s encoder-decoder architecture is applied to trajectory prediction—a numerical time-series prediction task—using a single value as the SOS can lead to a lack of smoothness in the transition between the first predicted value $\widehat{pos}_{\tau+1}$ and the observed trajectory $pos_{0:\tau}$. That is to say, $\widehat{pos}_{\tau+1}$ may deviate noticeably from the ground truth $pos_{\tau+1}$. This issue likely arises because, during the initial decoding stage, the Transformer decoder relies solely on the self-attention with SOS and the cross-attention with the encoder’s output. Insufficient information or excessive noise in the SOS may lead alignment bias, thus reducing prediction accuracy.
Figure 4: Comparison of SSOS and SOS.
To mitigate this issue, we propose using the sequence data $pos_{0,1,...,\tau}$ as the SOS. We refer to this approach as SSOS (Figure 4) and use it during training. In addition to computing the loss of the predicted trajectory $\widehat{pos}_{\tau+1:T-1}$, the loss of the reconstructed observed trajectory $\widehat{pos}_{1:\tau}$ is also minimized through gradient descent. The loss function using SSOS is presented below:
$\mathcal{L}_{\text{SSOS}}$ $=\lambda_{1}\cdot\text{MSE}(\widehat{pos}_{1:\tau},pos_{1:\tau})$
$\quad+\lambda_{2}\cdot\text{MSE}(\widehat{pos}_{\tau+1:T-1},pos_{\tau+1:T-1})$ (3)
where $\widehat{pos}_{1:\tau}$ denotes the reconstructed observed trajectory from time $1$ to $\tau$; $\widehat{pos}_{\tau+1:T-1}$ denotes the predicted trajectory excluding the sampled goal; $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that balance the contributions of the reconstruction loss and the prediction loss; MSE is the Mean Squared Error.
However, the SSOS strategy generates additional outputs $\widehat{pos}_{1:\tau}$, which are not part of the target prediction. These byproducts may introduce additional computational overhead.
4.4 Seq-CVAE: LSTM-Based CVAE
RNN (typically LSTM [11]) and Conditional Variational Autoencoder (CVAE) [26] have been widely applied in pedestrian trajectory prediction [15, 40, 43, 1]. We combine the two approaches and propose a generator, Seq-CVAE (Sequence CVAE), which is designed for sampling trajectory goals. The model predicts goals given the condition $c$, which is derived from the observed trajectory $pos_{0:\tau}$. These goals are subsequently fed into the Interaction Encoder to achieve its goal-driven functionality. Our purpose in this section is to learn a generative model $g_{\phi}(\widehat{pos}_{T}|c)$, where ${\phi}$ represents the model parameters. This model generates trajectory goals $\widehat{pos}_{T}$ that conform to the distribution of the conditional variable $c$. Seq-CVAE architecture is illustrated in Figure 5.
CVAE was first introduced in the field of image processing [26]. It compresses high-dimensional data, such as pixel-based images, into a low-dimensional latent space while preserving key features of the image. However, unlike images, trajectories typically consists of low-dimensional data, such as the coordinates in $pos_{0:\tau}$. To extract latent features like velocity, direction, and acceleration, these inputs require dimensionality expansion rather than compression. Consequently, we propose modifying the CVAE Encoder’s dimensionality reduction mechanism to perform dimensionality expansion instead.
Furthermore, regarding the construction of condition $c$ in CVAE, Yue et al. [43] extracts features from observed trajectory using an MLP. To better preserve the temporal feature of the trajectory, we adopt an LSTM module. To maintain information balance within the Encoder, the last step hidden state $h_{\tau}$ of the LSTM is passed through a fully connected layer for dimensionality reduction before concatenation, as shown in Figure 5. Therefore, for the Seq-CVAE, the condition $c$ for encoder and decoder is defined as
$c_{Encoder}$ $=\mathrm{\textit{FC}}(\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\}))$ (4)
$c_{Decoder}$ $=\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\})$ (5)
where $FC$ represents the fully connected layer, $h_{\tau}$ denotes the hidden state of LSTM at the last observed time step $\tau$, and $pos_{0:\tau}$ represents the coordinates of observed trajecotry.
[…]
5 Experiment
Datasets. We use the well-established datasets in the field of pedestrian trajectory prediction: ETH [22] and UCY [16]. The ETH/UCY datasets consist of five sub-datasets, covering four different scenarios and 1536 trajectories. These pedestrian trajectories include various interaction scenarios such as walking together, as couples, and in groups. Similar to previous work [31, 43, 42], we adopt the leave-one-out strategy for splitting the training and testing sets. That is, four datasets are used for training and validation, and one dataset is used for testing. Since partitioning the neighboring regions requires real-world distances, we convert the coordinate scale to real-world metrics in meters. Due to the different frame rates of the ETH and UCY datasets, we unify both to 2.5 fps, corresponding to a time step of 0.4 seconds.
Evaluation Protocol. Like many previous work [40, 41, 42], visual information is not used in our approach. During the test phase, our model takes the trajectory within 3.2 seconds as input and outputs the trajectory for the subsequent 4.8 seconds. For fair comparison, we adopt the best-of-$K$ protocol, which is sampling $K=20$ times and choose the best result for metric computation. This is the standard protocol in pedestrian trajectory prediction [17, 43, 7, 41].
Metrics. We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics. ADE refers to the average displacement error of the predicted trajectory $\widehat{pos}_{\tau+1:T}$ compared to the ground truth trajectory $pos_{\tau+1:T}$ and FDE represents the displacement error of the predicted goal $\widehat{pos}_{T}$ compared to the ground truth goal $pos_{T}$.
$ADE=\frac{1}{T-\tau}\sum_{t=\tau+1}^{T}\left\|\widehat{pos}_{t}-pos_{t}\right\|$ (6)
$FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (7)
where $\tau$ represents the last time step of the observed trajectory; $T$ represents the last time step of the predicted trajectory; $\widehat{pos}_{t}$ and $pos_{t}$ represent the predicted coordinates and ground truth coordinates at time step $t$, respectively.
Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE) metric, which will be discussed in detail in Section 5.2.
Implementation Details. In our model, the Transformer component adopts the original Transformer architecture [29]. For the Neighbor Encoder module, the hidden size is set to 128. In the Seq-CVAE module, the LSTM hidden size and the latent size are configured to 256, and the MLP parameter design is illustrated in Figure 5. In the data preprocessing stage, the neighbor radius $r$ for social interaction is set to 2 meters, which is determined based on the average ADE performance. Due to significant variations in the walking speed and direction across different scenes, this study employs data augmentation by rotatating trajectories at 0°, 90°, 180°, and 270°, along with scaling the trajectory length by a factor of 2. Notably, the direction-based interaction information $\mathcal{N}$ is transformed accordingly during rotation. To ensure the stability of the data distribution, all trajectories are shifted so that they have start at coordinate (0,0). For the distance features in social interaction information, the reciprocal of the distance is taken, meaning the diminishing influence of pedestrians farther away. These values are subsequently normalized using Min-Max normalization.
During training, the main components of InSyn (Interaction Encoder and Trajectory Generator) and the Seq-CVAE Goal Sampler are trained independently and the teacher forcing technique is applied to facilitate training. For the Encoder-Generator, the SSOS strategy is applied. The learning rate is set to $1\times 10^{-4}$. For the Seq-CVAE, the KL divergence weight in the loss function is configured to 5 to balance the model’s reconstruction capability and latent space distribution requirements, and the learning rate is set to $1\times 10^{-3}$. The model is optimized using the Adam optimizer with 50 epochs. Mean Squared Error (MSE) serves as the loss function. During testing, the model generates predictions autoregressively. All experiments are conducted on a single NVIDIA RTX 4090 GPU, hosted on the AutoDL cloud platform.
5.1 Comparison
Table 1: Results comparison on all datasets. Evaluation is performed using the Best-of-20 protocol strategy. The lower the better. Bold Represents the Best.
Methods Year ETH Hotel Univ Zara01 Zara02 Average
ADE FDE ADE FDE ADE FDE ADE FDE ADE FDE ADE FDE
SoPhie [24] 2019 0.70 1.43 0.76 1.67 0.54 1.24 0.30 0.63 0.38 0.78 0.54 1.15
CGNS [17] 2019 0.62 1.40 0.70 0.93 0.48 1.22 0.32 0.59 0.35 0.71 0.49 0.97
TF [9] 2021 0.61 1.12 0.18 0.30 0.35 0.65 0.22 0.38 0.17 0.32 0.31 0.55
PECNet [7] 2021 0.54 0.87 0.18 0.24 0.35 0.60 0.22 0.39 0.17 0.30 0.29 0.48
SimFuse [10] 2021 0.59 1.18 0.31 0.73 0.50 1.17 0.27 0.54 0.27 0.56 0.38 0.84
STG-DAT[18] 2021 0.38 0.77 0.25 0.39 0.41 0.82 0.23 0.50 0.21 0.46 0.30 0.59
TDAGCN [32] 2023 0.52 0.72 0.26 0.46 0.32 0.53 0.26 0.45 0.18 0.29 0.31 0.49
MSTCNN [25] 2024 0.63 0.98 0.32 0.49 0.42 0.72 0.32 0.50 0.28 0.44 0.39 0.63
Goal-CurveNet [33] 2024 0.45 0.68 0.36 0.68 0.31 0.48 0.34 0.58 0.23 0.45 0.34 0.57
MSWTE-GNN [19] 2025 0.51 1.04 0.23 0.44 0.32 0.64 0.24 0.45 0.23 0.42 0.31 0.60
InSyn (Ours) - 0.36 0.77 0.27 0.47 0.31 0.54 0.20 0.44 0.15 0.36 0.26 0.52
5.1.1 Comparison Baselines
We compare our model with representative methods PECNet [7] and TF [9], as well as the latest work [33, 32, 25, 19]. Additionally, we include some common baselines such as SoPhie [24] and CGNS [17]. The baseline results are from officially-reported metrics.
Among these, SoPhie [24], CGNS [17], STG-DAT [18], and Goal-CurveNet [33] leverage visual information, which our approach does not utilize. TF [9] is the first to apply Transformers for human trajectory prediction; however, unlike our method, it focuses on individual motion without explicitly modeling social or environmental interactions.
5.1.2 Evaluation
The comparative results are demonstrated in Table  1. Compared to the aforementioned methods, our model achieves the best average ADE of 0.26. For goal prediction, the Seq-CVAE achieves excellent performance with an average FDE of 0.52, slightly trailing only behind PECNet [7] and TDAGCN [32] despite its relatively lightweight design. In contrast to FDE, which only evaluates the final predicted position, ADE measures error across all predicted trajectory points, with greater emphasis on intermediate trajectory deviations, thereby better reflecting the model’s capacity to model complex pedestrian interactions. To focus on the model’s ability to capture interactions, we adopt ADE as the evaluation metric for assessing and comparing performance across different scenarios.
Crowded Scenarios. In the ETH subset, which represents a bidirectional passageway with frequent pedestrian interactions, InSyn accurately captures complex behaviors such as walking In Sync and Conflict, and demonstrates strong overall results in this challenging scenario. Similarly, in the UCY dataset, where human-human interactions dominate, InSyn excels on the Zara01 and Zara02 subsets, achieving leading ADE performance. On the Univ subset, our model’s performance is comparable to Goal-CurveNet [33], which leverages scene visual information, exhibiting equivalent top-tier results.
Sparse Scenarios. However, on the Hotel dataset, our model underperforms relatively compared to original Transformer [9]. The Hotel dataset is a smaller scene area, which limits the effectiveness of InSyn’s distance-based interaction region partitioning. Moreover, the dataset contains predominantly linear trajectories with sparse interactions. In this scenario, the basic attention mechanism of the original Transformer is sufficient to capture motion patterns effectively, and the additional complex interaction modeling in our model may lead to overfitting, causing excessive reliance on interaction behaviors. This issue is also shared by other complex models such as [32] and [33], as shown in Table 1.
Overall, our model achieves the best average ADE among the compared models and performs particularly well in scenarios with more complex interactions. This demonstrates the model’s ability to effectively handle intricate interactions among pedestrians.
5.2 Ablation Studies
5.2.1 Interaction-Related Components Analysis
For the interaction part, region partition divides the circular area centered on the agent’s position into four regions to capture directional interaction effects. The interaction state, proposed in this study, categorizes interactions into three types: In Sync, Conflict, and No Neighbor. To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table 2. These variations allow us to isolate the individual and combined effects of region partition and interaction state. The original InSyn incorporates both region partition and interaction state. Case studies on the variations are provided in the supplementary material.
From the results in Table  2, it is evident that the model incorporating both region rartition and interaction state, namely the original InSyn, achieves the best performance. This demonstrates that explicitly dividing interaction behaviors enhances the model’s ability to extract complex interaction patterns among an agent’s neighbors, leading to more reasonable trajectory predictions.
Specifically, in the absence of region partition, the model may struggle to discern the direction of influence, relying solely on training data distributions to infer future trajectory deviations—a limitation that undermines its generalizability. Similarly, if without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance. This would underperform when the training and testing sets differ. For instance, if the training set contains more In Sync scenarios, the trained model may underestimate repulsive effects in Conflict situations.
Table 2: ADE Result of Ablation Study on Region Partition and Interaction State. RP: region partition; IS: interaction state. w/o denotes Without. InSyn incorporates both region partition and interaction state, whereas the Baseline model includes neither.
Methods ADE
ETH Hotel Univ Zara01 Zara02 Average
InSyn 0.36 0.27 0.31 0.20 0.15 0.26
w/o-RP 0.61 0.33 0.37 0.28 0.23 0.36
w/o-IS 0.46 0.27 0.36 0.25 0.29 0.33
Baseline 0.60 0.28 0.41 0.29 0.24 0.36
5.2.2 SSOS Strategy Analysis
Additionally, we conducted an ablation study on the SSOS strategy proposed in this paper. SSOS is designed to mitigate the divergence at the first step when using Transformer’s Encoder-Decoder architecture for numerical time-series prediction tasks, such as pedestrian trajectory prediction. In this context, steps $0:\tau$ represent observed input to the encoder, while time steps $\tau+1:T$ correspond to the future trajectory to be predicted. We replace the traditional single SOS with the sequence-based SSOS. This approach incorporates the observed values from steps $0:\tau$ as inputs for both the encoder and decoder during the initial stage. During the autoregressive process, the outputs are progressively concatenated and fed back into the Trajectory Generator as inputs. This allows the model to predict the complete trajectory while computing the loss for all outputs (Equation 3). To evaluate the effectiveness of SSOS, we introduce the Initial Displacement Error (IDE):
$IDE=\left\|\widehat{pos}_{\tau+1}-pos_{\tau+1}\right\|$ (8)
where $\tau+1$ denotes the first prediction time step.
Table  3 compares the average IDE results across all datasets using SSOS and SOS (set to $pos_{\tau}$). The results show that the SSOS strategy reduces the average IDE by approximately 6.58% compared to the SOS strategy. This reduction alleviates prediction divergence at time $\tau+1$, enabling smoother transitions between observed and predicted trajectories. Additionally, it mitigates error accumulation during autoregressive prediction.
Table 3: IDE Result of Ablation Study on SSOS. InSyn-SSOS: trained using SSOS; InSyn-SOS: trained using SOS. InSyn-SSOS represents the original InSyn model.
Methods IDE
ETH Hotel Univ Zara01 Zara02 Average
InSyn-SSOS 0.116 0.054 0.083 0.051 0.053 0.071
InSyn-SOS 0.118 0.051 0.085 0.062 0.065 0.076
5.3 Visualization
To further explore its practical performance, we conducted a qualitative analysis of the model’s behavior in real-world scenarios via visualization. Figure 6 illustrates the walking state of a specific agent (pedestrian 1) during the observation time steps, with details on its trajectory and interactions. After feeding this input into InSyn, the model outputs the predicted trajectory distribution, which is visualized using Kernel Density Estimation (KDE).
In this example, pedestrian 2 moves in the same direction as the agent, suggesting synchronization. With the exception of a brief Conflict at $t=3$, when pedestrian 2 temporarily moves to the upper-right of the agent, pedestrian 2 remains largely In Sync with the agent. Consequently, despite consistently staying to the agent’s vicinity, pedestrian 2 exerts relatively weak influence on the agent’s trajectory. At $t=7$, pedestrian 3 enters the agent’s neighbor region while moving in the opposite direction. The model encodes this information—"a Conflict from the upper-left region at $t=7$, at a distance of $d_{t=7}^{lu}$"—into a representation. The LSTM aggregates representations across history $0:\tau$ ($\tau=7$) and the Trajectory Generator interprets its hidden state $h_{\tau}$ as an "avoidance" signal. As a result, during the initial prediction steps, the agent turns to the right to avoid pedestrian 3. Since the agent consistently walks toward the upper-right during the observation period, the goals sampled by Seq-CVAE are predominantly directed toward the upper-right. After predicting the "avoidance" behavior, the goal-driven attention mechanism guides the model to refocus on moving toward the sampled goal.
Due to the inherent randomness in pedestrian trajectories (e.g., variations in speed), long-term goal predictions may occasionally deviate from the ground truth. However, intermediate trajectory predictions, including directional changes, remain relatively accurate (see Figure 6). Through visualizing pedestrian trajectories, the model demonstrates its sensitivity to interaction relationships and its ability to effectively leverage goal-driven predictions. More case studies are provided in the supplementary material.
Figure 6: Visualization of Predicted Trajectory Distributions. We visualized the trajectory prediction of an agent (pedestrian 1) in a scene where In Sync, Conflict, and No Neighbor occur. The number of samples is $K$=20, meaning 20 goals were sampled. For each goal, a trajectory was generated in a goal-driven manner. Since pedestrian 4 remains outside the agent’s neighbor region, the model disregards its influence.
6 Conclusion
This paper introduces InSyn, a novel model that extends the Transformer architecture to explicitly capture complex pedestrian interaction patterns. Experimental results demonstrate that our model excels in scenarios involving intricate interaction behaviors, effectively modeling detailed temporal and spatial relationships. Furthermore, to mitigate initial prediction divergence when applying the Transformer’s encoder-decoder framework to numerical sequence data, we propose the SSOS strategy, which relatively alleviates this challenge.
Future work will focus on two key directions. First, we aim to design a more refined approach for partitioning interaction regions to better handle challenges posed by small-scale scenes. Second, we plan to apply SSOS strategy to other types of numerical sequential data beyond trajectory prediction, exploring its potential in diverse domains.
{ack}

### v2

@@
-InSyn: Modeling Complex Interactions for Pedestrian Trajectory Prediction
+Trustworthy Pedestrian Trajectory Prediction via Pattern-Aware Interaction Modeling
 Abstract
-Accurate pedestrian trajectory prediction is crucial for intelligent applications, yet it remains highly challenging due to the complexity of interactions among pedestrians.
-Previous methods have primarily relied on relative positions to model pedestrian interactions; however, they tend to overlook specific interaction patterns such as paired walking or conflicting behaviors, limiting the prediction accuracy in crowded scenarios.
+Accurate and reliable pedestrian trajectory prediction is critical for the safety and robustness of intelligent applications, yet achieving trustworthy prediction remains highly challenging due to the complexity of interactions among pedestrians.
+Previous methods often adopt black-box modeling of pedestrian interactions, treating all neighbors uniformly.
+Despite their strong performance, such opaque modeling limits the reliability of predictions in safety-critical real-world deployments.
 To address this issue, we propose InSyn (Interaction-Synchronization Network), a novel Transformer-based model that explicitly captures diverse interaction patterns (e.g., walking in sync or conflicting) while effectively modeling direction-sensitive social behaviors.
-Additionally, we introduce a training strategy termed Seq-Start of Seq (SSOS), designed to alleviate the common issue of initial-step divergence in numerical time-series prediction.
-Experiments on the ETH and UCY datasets demonstrate that our model outperforms recent baselines significantly, especially in high-density scenarios.
-Furthermore, the SSOS strategy proves effective in improving sequential prediction performance, reducing the initial-step prediction error by approximately 6.58%.
-[…]
-7556
-1 Introduction
-Pedestrian trajectory prediction plays a crucial role in various fields such as autonomous driving [21], robotic navigation [4], and smart cities [5].
-Accurately forecasting a pedestrian’s future trajectory not only helps intelligent systems better understand human behavior but also significantly enhances the safety and efficiency of human-machine interaction.
-However, beyond an individual’s own motion, interactions among pedestrians are often highly complex, making trajectory prediction a challenging task.
+Additionally, we introduce a training strategy, termed Seq-Start of Seq (SSOS), designed to alleviate the common issue of initial-step divergence in numerical time-series prediction.
+Experiments on the ETH and UCY datasets demonstrate that our model not only outperforms recent black-box baselines in prediction accuracy, especially under high-density scenarios, but also provides stronger interpretability, achieving a favorable trade-off between reliability and accuracy.
+Furthermore, the SSOS strategy proves to be effective in improving sequential prediction performance, reducing the initial-step prediction error by approximately 6.58%.
+Index Terms:
+Human Trajectory Prediction, Interpretable Modeling, Attention Mechanism
+I Introduction
+Pedestrian trajectory prediction is essential for safety-critical applications such as autonomous driving [1] and robotic navigation [2].
+Accurate and reliable forecasting enables intelligent systems to understand human behavior better and ensures safe and trustworthy predictions.
+However, the presence of complex pedestrian interactions poses significant challenges to the task.
 In a given scene, pedestrians may respond to nearby individuals in different ways.
@@
 Alternatively, they may show weak or no influence on each other, such as when walking in sync or as part of a group.
-Notably, even in walk-in-sync scenarios, subtle interactions may still occur.
+In particular, even in walk-in-sync scenarios, subtle interactions may still occur.
 Effectively modeling these intricate interaction patterns, especially in high-density environments, remains a major challenge in pedestrian trajectory prediction.
@@
 Our Approach.
-F represents the interaction affect between pedestrians.
+F represents the interaction effect between pedestrians.
 In traditional approaches (top), all neighbors of the agent are treated as being in the same state, with interactions solely based on the distance between the agent and its neighbors.
@@
 Recently, extensive studies have investigated pedestrian interactions [3, 4, 5, 6], using approaches like social pooling layers [3, 7], social force mechanism [6], and graph neural networks (GNNs) [8, 9].
-However, a common characteristic of these methods is that they rely on relative positions between an agent and its neighbors to model influence, without considering the specific interaction relationships between them.
-This limitation restricts the capacity of previous models to capture the dynamics of real-world interactions, such as walking in sync or avoiding collisions, which are crucial for accurately predicting trajectory in crowded scenario.
-Prior approaches treat all neighboring pedestrians uniformly, failing to differentiate patterns such as In Sync and Conflict (see Figure 1).
-This oversimplification may leads to overfitting on sparse datasets and underperforming in complex social scenarios.
-For instance, in situations where pedestrians walk in sync (e.g., friends walking side-by-side), failing to account for this coordinated behavior may lead to overlapping trajectories during prediction.
-Similarly, in conflict scenarios (e.g., two pedestrians approaching each other head-on or navigating narrow passages), the lack of explicit modeling for competitive or avoidance behaviors often results in unrealistic trajectories that either overestimate collision risk or underestimate the necessity for trajectory alteration.
+However, a common characteristic of these methods is that they rely on relative positions between an agent and its neighbors to model influence, treat all neighboring pedestrians by black-box representation, failing to differentiate patterns such as In Sync and Conflict (see Figure 1).
+This oversimplification may lead to overfitting and untrustworthy predictions.
+For example, ignoring synchronized walking (e.g., friends walking side-by-side) can lead to overlapping predicted paths.
+In conflict cases, such as head-on encounters, the absence of explicit modeling may cause unrealistic predictions that either overestimate collisions or miss necessary evasive actions.
+This black-box modeling not only reduces prediction accuracy, but also limits interpretability and trustworthiness in safety-critical scenarios such as machine navigation and intelligent transportation systems.
 These limitations become particularly pronounced in crowded or dynamic environments.
-To overcome this issue, we need a method that explicitly identifies and captures these interaction patterns, enabling a deeper understanding of the diverse pedestrian interactions, as shown in Figure 1.
+To address this issue, we need a method that understands what’s exactly going on between pedestrians—recognizing concrete interaction patterns rather than treating them as abstract black-box inputs—to ensure more reliable and interpretable predictions (see Figure 1).
 Although pedestrian trajectories are highly stochastic, they usually have a clear goal [10].
@@
 Experimental results from Giuliari et al.
-[9] demonstrate that the attention mechanism is well-suited for processing sequential data with irregular time intervals (e.g., observed trajectories and goal), effectively capturing long-term dependencies in pedestrian motion and providing reliable goal-driven functionality.
+[11] demonstrate that the attention mechanism is well-suited for processing sequential data with irregular time intervals (e.g., observed trajectories and goals), effectively capturing long-term dependencies in pedestrian motion and providing reliable goal-driven functionality.
 This paper proposes the InSyn, a Transformer-based model.
@@
 (2) Trajectory Generator: forms an attention mechanism in conjunction with the Interaction Encoder, and incorporates the proposed SSOS strategy to alleviate divergence at the initial prediction step.
-(3) Seq-CVAE Goal Sampler: a conditional generative model specifically designed for sequential prediction, used for goals sampling.
+(3) Seq-CVAE Goal Sampler: a conditional generative model specifically designed for sequential prediction, used for goal sampling.
 We evaluate our model on the well-established pedestrian trajectory prediction datasets ETH [12] and UCY [13].
 The experimental results demonstrate that our model outperforms recent methods in key metrics, particularly in scenarios with complex interactions.
-We further conduct ablation studies to show the superiority of interaction modeling and SSOS strategy.
+We further conduct ablation studies to show the superiority of interaction modeling and the SSOS strategy.
+Additionally, the case study highlights the interpretability of our approach by illustrating how specific interaction patterns influence prediction outcomes.
 Figure 2: Left is the illustration of the input $S_{0:\tau}$.
@@
 In summary, our work makes two key contributions:
-  • We propose a pedestrian interaction modeling approach that explicitly identifies and leverages specific interaction patterns, achieving a significant improvement in average ADE compared to the previous baselines and offering more comprehensive insights into socially-aware motion patterns.
+  • We propose a pedestrian interaction modeling approach that explicitly identifies and leverages specific interaction patterns, achieving a significant improvement in average ADE compared to the previous black-box baselines and contributing to more trustworthy socially-aware trajectory prediction.
   • For numerical time-series prediction tasks, we introduce a novel training strategy for the Transformer encoder-decoder architecture.
 This strategy, termed SSOS, mitigates the divergence in the first prediction step, thereby reducing error accumulation and improving overall performance.
-2 Related Work
-2.1 Pedestrian Trajectory Prediction
+II Related Work
+II-A Pedestrian Trajectory Prediction
 Pedestrian trajectory prediction methods can be categorized into traditional machine learning approaches and deep learning-based approaches.
@@
 [21] proposed a variant LSTM structure for information sharing within scenes, while SR-LSTM [22] combined neighbor intentions and message passing to enhance prediction accuracy.
-These work leverage the strong temporal feature extraction capabilities of LSTM.
-In recent years, Transformer have gained popularity in trajectory prediction due to their ability to capture long-range dependencies.
+These works leverage the strong temporal feature extraction capabilities of LSTM.
+In recent years, Transformers have gained popularity in trajectory prediction due to their ability to capture long-range dependencies.
 Liu et al.
-[20] combined CNN and Transformer, using attention mechanism to process interaction information.
+[23] combined CNN and Transformer, using the attention mechanism to process interaction information.
 MRGTraj [24] integrated a Transformer encoder with a novel decoder, achieving trajectory prediction through a mapping-refining-generating structure.
@@
 Building on these advancements, our research combines the strengths of LSTM and Transformer.
-By leveraging the temporal feature extraction capabilities of LSTM, we enhance the Transformer’s attention mechanism with processed temporal interactions features, thereby improving prediction performance.
-2.2 Pedestrian Interaction Modeling
+By leveraging the temporal feature extraction capabilities of LSTM, we enhance the Transformer’s attention mechanism with processed temporal interaction features, thereby improving prediction performance.
+II-B Pedestrian Interaction Modeling
 Pedestrian interaction is a key factor influencing prediction accuracy. Pedestrian motion is not only driven by individual goals but also affected by interactions with surrounding pedestrians.
@@
 [28] introduced a social graph attention mechanism combined with a pseudo-oracle predictor to capture social interactions and intention states, enhancing prediction accuracy.
-Transformer have been utilized in pedestrian interaction modeling [34, 39, 12, 2, 27].
+Transformers have been utilized in pedestrian interaction modeling [29, 30, 31, 32, 33].
 For example, Yuan et al.
@@
 Similarly, Yang et al.
-[39] leveraged GNNs and Transformer to model spatial and temporal dependencies.
+[30] leveraged GNNs and Transformers to model spatial and temporal dependencies.
 To address future interaction modeling, Amirloo et al.
 [32] employed the self-attention mechanism of Transformer to model pedestrian interactions and considered future states autoregressively during decoding to avoid trajectory conflicts.
-However, most of these methods fail to consider the specific interaction relationships among pedestrians, such as walking In Sync or engaging in Conflict.
-To address this limitation, we propose a model that explicitly captures specific types of interaction features.
-3 Problem Formulation
+However, most of these methods rely on black-box representations of social influence.
+This limits the model’s ability to produce reliable and interpretable predictions, especially in complex or safety-critical scenarios.
+In contrast, our approach explicitly captures structured interaction patterns, such as walking In Sync or engaging in Conflict, aiming to enhance the trustworthiness and transparency of socially-aware trajectory prediction.
+III Problem Formulation
 In this study, pedestrian trajectory prediction is formulated as follows: given the walking states $S_{0:\tau}$ of a certain agent during the observation time steps $\mathord{0:\tau}$, the model outputs the predicted trajectory $\widehat{pos}_{\tau+1:T}$ of the agent for the future time steps $\mathord{\tau+1:T}$.
@@
 $N_{t}=\left[\left(f_{t}^{lu},d_{t}^{lu}\right),\left(f_{t}^{ru},d_{t}^{ru}\right),\left(f_{t}^{ld},d_{t}^{ld}\right),\left(f_{t}^{rd},d_{t}^{rd}\right)\right]$ (1)
-where $f_{t}^{lu},f_{t}^{ru},f_{t}^{ld},f_{t}^{rd}$ represent the interaction states in the left-up, right-up, left-down, and right-down regions respectively.
+where $f_{t}^{lu},f_{t}^{ru},f_{t}^{ld},f_{t}^{rd}$ represent the interaction states in the left-up, right-up, left-down, and right-down regions, respectively.
 These patterns include No Neighbor, In Sync, and Conflict, as shown in Figure 2.
-Specifically, if the nearest neighbor in a given region at time $t$ is the same as the one at time $t-1$, the interaction state is classified as In Sync; otherwise, it is marked as Conflict.
+The construction of interaction patterns is highly transparent and trustworthy, as it relies on interpretable spatial-temporal neighbor changes.
+Specifically, if the nearest neighbor in a given region at time $t$ is the same as the one at time t1t-1, the interaction state is classified as In Sync; otherwise, it is marked as Conflict.
 In cases of No Neighbor or In Sync, the interaction influence of the neighbor on the agent’s future trajectory may be negligible.
-However, if a new pedestrian suddenly enters in a region, referred as Conflict, it is more likely to impact the agent’s trajectory.
+However, if a new pedestrian suddenly enters a region, referred to as Conflict, it is more likely to impact the agent’s trajectory.
 The terms $d_{t}^{lu},d_{t}^{ru},d_{t}^{ld},d_{t}^{rd}$ indicate the distances to the nearest neighbor in each region.
@@
 where $\theta$ represents the model parameters.
-The model aims to predict the distribution of pedestrian’s future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
+The model aims to predict the distribution of pedestrians’ future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
 Figure 3: Overview of the InSyn framework for trajectory prediction.
 Our model consists of three key modules: (1) Interaction Encoder, (2) Trajectory Generator, and (3) Seq-CVAE.
-The input observed walking state includes the agent’s trajectory positions $pos_{0:\tau}$ and its interaction information $N_{0:\tau}$ within the obserbed time $0:\tau$.
-4 Methodology
-4.1 Framework Design
-In this section, we propose InSyn, a Transformer-based model designed to identify and model spefific social interaction.
+The input observed walking state includes the agent’s trajectory positions $pos_{0:\tau}$ and its interaction information $N_{0:\tau}$ within the observed time $0:\tau$.
+IV Methodology
+In this section, we propose InSyn, a Transformer-based model tailored for the interpretable modeling of specific social interactions.
 The model consists of three components: (1) Interaction Encoder: Extracts interaction features and enables goal-driven prediction.
-(2) Trajectory Generator: Employing SSOS strategy to mitigate initial prediction step divergence.
+(2) Trajectory Generator: Employing the SSOS strategy to mitigate initial prediction step divergence.
 (3) Seq-CVAE Goal Sampler: A generator for sequential data that utilizes high-dimensional latent variables for goal sampling.
@@
 We then elaborate on the motivation behind the SSOS strategy and, finally, the design of the Seq-CVAE module.
-4.2 Interaction Encoder
+IV-A Interaction Encoder
 The Interaction Encoder integrates observed trajectory, interaction information, and sampled goal to extract spatiotemporal features.
 These features are utilized for goal-driven and interaction-sensitive trajectory prediction.
-4.2.1 Goal-Driven
-Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  4.3) is combined with the position information through concatenation.
+IV-A1 Goal-Driven
+Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  IV-B) is combined with the position information through concatenation.
 Notably, since self-attention cannot inherently capture positional information, positional encoding is used to encode the input positions.
@@
 Therefore, before concatenating $pos_{0:\tau}^{p}$ with the sampled goal $\widehat{pos}_{T}$, padding needs to be inserted between them to ensure that the goal is assigned the corresponding positional encoding.
-4.2.2 Interaction
-The complex interactions among pedestrians are extracted through a neighbor encoder module follwed by an LSTM.
+IV-A2 Interaction
+To facilitate trustworthy modeling of social interactions, we introduce a modular neighbor encoder that encodes pattern-aware and distance-based interaction effects.
+The interaction patterns between pedestrians are extracted through a neighbor encoder module followed by an LSTM.
 The agent $p$’s social interaction information, represented as $\mathcal{N}^{p}=\left(N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}\right)$, are encoded using the neighbor encoder.
-The design of the neighbor encoder includes three components: the embedding layer, the Distance Gate, and the Social Pooling Layer.
-The embedding layer maps three interaction states (No Neighbor, In Sync, Conflict) into corresponding representations.
+which consists of three components: an embedding layer, a Distance Gate, and a Social Pooling Layer.
+The embedding layer maps interaction states (No Neighbor, In Sync, Conflict) to learnable representations.
 These representations are then combined with the processed distance $d$ through element-wise multiplication.
 The process of $d$ is through the Distance Gate: $\mathrm{Gate}(d)=\sigma(\mathbf{W}\cdot d+\mathbf{b})$, where $\sigma(z)=\frac{1}{1+e^{-z}}$ and $\mathbf{W},\mathbf{b}$ represent the parameters of linear layer.
-The Distance Gate is similar to the gate mechanism in LSTM [11].
-It maps $d$ to a weight in the range [0,1].
-This weight reflects distance-based intensity, enhancing the spatial characteristics of pedestrian interactions.
+Inspired by the gating mechanism in LSTM [20], the Distance Gate maps $d$ to a weight in [0,1], reflecting distance-based interaction intensity.
 After forming four regional interaction representations, they are passed through the Social Pooling Layer, which applies a max-pooling function across the region dimension.
@@
 Consequently, $h_{\tau}$ is directly concatenated with the positionally encoded representation (see Figure 3).
-4.3 SSOS: Seq-Start of Seq
-The Trajectory Generator incorporates the SSOS Strategy into the Transformer Decoder [29] to alleviate initial prediction divergence.
-The Transformer’s decoder relies on the Start of Sequence (SOS) as the initial input when generating the first output.
-In Natural Language Processing (NLP), the Begin-of-Sentence token $bos\right>$ is commonly used as the SOS, which aligns with the semantic structure of text.
-However, in sequential trajectory prediction, the selection of SOS should consider the spatial information; otherwise, it may introduce noise.
-Giuliari et al.
-[9] set the SOS to $(0,0)$ for trajectories prediction.
-While $(0,0)$ represents a fixed position in spatial coordinates, using it as the SOS may introduce noise and mislead the model due to its lack of alignment with the observed trajectory.
-Therefore, it is reasonable to consider using the position at the last observed time step ,$\widehat{pos}_{\tau}$, as the SOS token of decoder.
+IV-B SSOS: Seq-Start of Seq
+[…]
 Nevertheless, we note that when the Transformer’s encoder-decoder architecture is applied to trajectory prediction—a numerical time-series prediction task—using a single value as the SOS can lead to a lack of smoothness in the transition between the first predicted value $\widehat{pos}_{\tau+1}$ and the observed trajectory $pos_{0:\tau}$.
 That is to say, $\widehat{pos}_{\tau+1}$ may deviate noticeably from the ground truth $pos_{\tau+1}$.
-This issue likely arises because, during the initial decoding stage, the Transformer decoder relies solely on the self-attention with SOS and the cross-attention with the encoder’s output.
-Insufficient information or excessive noise in the SOS may lead alignment bias, thus reducing prediction accuracy.
-Figure 4: Comparison of SSOS and SOS.
+This issue likely arises because, during the initial decoding stage, the Transformer decoder relies solely on the self-attention with the SOS and the cross-attention with the encoder’s output.
+Insufficient information or excessive noise in the SOS may lead to alignment bias, thus reducing prediction accuracy.
 To mitigate this issue, we propose using the sequence data $pos_{0,1,...,\tau}$ as the SOS.
-We refer to this approach as SSOS (Figure 4) and use it during training.
+We refer to this approach as SSOS and use it during training.
 In addition to computing the loss of the predicted trajectory $\widehat{pos}_{\tau+1:T-1}$, the loss of the reconstructed observed trajectory $\widehat{pos}_{1:\tau}$ is also minimized through gradient descent.
@@
 However, the SSOS strategy generates additional outputs $\widehat{pos}_{1:\tau}$, which are not part of the target prediction. These byproducts may introduce additional computational overhead.
-4.4 Seq-CVAE: LSTM-Based CVAE
+IV-C Seq-CVAE: LSTM-Based CVAE
 RNN (typically LSTM [20]) and Conditional Variational Autoencoder (CVAE) [36] have been widely applied in pedestrian trajectory prediction [37, 21, 6, 3].
@@
 This model generates trajectory goals $\widehat{pos}_{T}$ that conform to the distribution of the conditional variable $c$.
-Seq-CVAE architecture is illustrated in Figure 5.
+The Seq-CVAE architecture is illustrated in Figure 4.
 CVAE was first introduced in the field of image processing [36].
 It compresses high-dimensional data, such as pixel-based images, into a low-dimensional latent space while preserving key features of the image.
-However, unlike images, trajectories typically consists of low-dimensional data, such as the coordinates in $pos_{0:\tau}$.
+However, unlike images, trajectories typically consist of low-dimensional data, such as the coordinates in $pos_{0:\tau}$.
 To extract latent features like velocity, direction, and acceleration, these inputs require dimensionality expansion rather than compression.
@@
 Furthermore, regarding the construction of condition $c$ in CVAE, Yue et al.
-[43] extracts features from observed trajectory using an MLP.
+[6] extract features from observed trajectory using an MLP.
 To better preserve the temporal feature of the trajectory, we adopt an LSTM module.
 To maintain information balance within the Encoder, the last step hidden state $h_{\tau}$ of the LSTM is passed through a fully connected layer for dimensionality reduction before concatenation, as shown in Figure 4.
-Therefore, for the Seq-CVAE, the condition $c$ for encoder and decoder is defined as
+Therefore, for the Seq-CVAE, the condition $c$ for the encoder and decoder is defined as
 $c_{Encoder}$ $=\mathrm{\textit{FC}}(\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\}))$ (4)
 $c_{Decoder}$ $=\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\})$ (5)
-where $FC$ represents the fully connected layer, $h_{\tau}$ denotes the hidden state of LSTM at the last observed time step $\tau$, and $pos_{0:\tau}$ represents the coordinates of observed trajecotry.
+where $FC$ represents the fully connected layer, $h_{\tau}$ denotes the hidden state of LSTM at the last observed time step $\tau$, and $pos_{0:\tau}$ represents the coordinates of the observed trajectory.
 […]
-5 Experiment
-Datasets.
-We use the well-established datasets in the field of pedestrian trajectory prediction: ETH [22] and UCY [16].
-The ETH/UCY datasets consist of five sub-datasets, covering four different scenarios and 1536 trajectories.
-These pedestrian trajectories include various interaction scenarios such as walking together, as couples, and in groups.
-Similar to previous work [31, 43, 42], we adopt the leave-one-out strategy for splitting the training and testing sets.
-That is, four datasets are used for training and validation, and one dataset is used for testing.
-Since partitioning the neighboring regions requires real-world distances, we convert the coordinate scale to real-world metrics in meters.
-Due to the different frame rates of the ETH and UCY datasets, we unify both to 2.5 fps, corresponding to a time step of 0.4 seconds.
-Evaluation Protocol.
-Like many previous work [40, 41, 42], visual information is not used in our approach.
-During the test phase, our model takes the trajectory within 3.2 seconds as input and outputs the trajectory for the subsequent 4.8 seconds.
-For fair comparison, we adopt the best-of-$K$ protocol, which is sampling $K=20$ times and choose the best result for metric computation.
-This is the standard protocol in pedestrian trajectory prediction [17, 43, 7, 41].
-Metrics.
-We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics.
-ADE refers to the average displacement error of the predicted trajectory $\widehat{pos}_{\tau+1:T}$ compared to the ground truth trajectory $pos_{\tau+1:T}$ and FDE represents the displacement error of the predicted goal $\widehat{pos}_{T}$ compared to the ground truth goal $pos_{T}$.
-$ADE=\frac{1}{T-\tau}\sum_{t=\tau+1}^{T}\left\|\widehat{pos}_{t}-pos_{t}\right\|$ (6)
-$FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (7)
-where $\tau$ represents the last time step of the observed trajectory; $T$ represents the last time step of the predicted trajectory; $\widehat{pos}_{t}$ and $pos_{t}$ represent the predicted coordinates and ground truth coordinates at time step $t$, respectively.
-Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE) metric, which will be discussed in detail in Section 5.2.
-Implementation Details.
-In our model, the Transformer component adopts the original Transformer architecture [29].
-For the Neighbor Encoder module, the hidden size is set to 128.
-In the Seq-CVAE module, the LSTM hidden size and the latent size are configured to 256, and the MLP parameter design is illustrated in Figure 5.
-In the data preprocessing stage, the neighbor radius $r$ for social interaction is set to 2 meters, which is determined based on the average ADE performance.
-Due to significant variations in the walking speed and direction across different scenes, this study employs data augmentation by rotatating trajectories at 0°, 90°, 180°, and 270°, along with scaling the trajectory length by a factor of 2.
-Notably, the direction-based interaction information $\mathcal{N}$ is transformed accordingly during rotation.
-To ensure the stability of the data distribution, all trajectories are shifted so that they have start at coordinate (0,0).
-For the distance features in social interaction information, the reciprocal of the distance is taken, meaning the diminishing influence of pedestrians farther away.
-These values are subsequently normalized using Min-Max normalization.
-During training, the main components of InSyn (Interaction Encoder and Trajectory Generator) and the Seq-CVAE Goal Sampler are trained independently and the teacher forcing technique is applied to facilitate training.
-For the Encoder-Generator, the SSOS strategy is applied.
-The learning rate is set to $1\times 10^{-4}$.
-For the Seq-CVAE, the KL divergence weight in the loss function is configured to 5 to balance the model’s reconstruction capability and latent space distribution requirements, and the learning rate is set to $1\times 10^{-3}$.
-The model is optimized using the Adam optimizer with 50 epochs.
-Mean Squared Error (MSE) serves as the loss function.
-During testing, the model generates predictions autoregressively.
-All experiments are conducted on a single NVIDIA RTX 4090 GPU, hosted on the AutoDL cloud platform.
-5.1 Comparison
-Table 1: Results comparison on all datasets. Evaluation is performed using the Best-of-20 protocol strategy. The lower the better. Bold Represents the Best.
+V Experiment
+TABLE I: Results comparison on all datasets. Bold: Best
 Methods Year ETH Hotel Univ Zara01 Zara02 Average
@@
 InSyn (Ours) - 0.36 0.77 0.27 0.47 0.31 0.54 0.20 0.44 0.15 0.36 0.26 0.52
-5.1.1 Comparison Baselines
+Datasets.
+We use the well-established datasets in the field of pedestrian trajectory prediction: ETH [12] and UCY [13].
+The datasets include five sub-datasets, covering four different scenarios and 1536 trajectories.
+These trajectories include various interaction scenarios such as walking together, as couples, and in groups.
+Similar to previous work [43, 6, 34], we adopt the leave-one-out strategy for splitting the training and testing sets.
+That is, four datasets are used for training and validation, and one dataset is used for testing.
+Since partitioning the neighboring regions requires real-world distances, we convert the coordinate scale to real-world metrics in meters.
+Due to the different frame rates of the ETH and UCY datasets, we unify both to 2.5 fps, corresponding to a time step of 0.4 seconds.
+Evaluation Protocol.
+Like many previous work [21, 8, 34], visual information is not used in our approach.
+During the test phase, our model takes the trajectory within 3.2 seconds as input and outputs the trajectory for the subsequent 4.8 seconds.
+For fair comparison, we adopt the best-of-$K$ protocol, which is sampling $K=20$ times and choosing the best result for metric computation.
+This is the standard protocol in pedestrian trajectory prediction [40, 6, 7, 8].
+Metrics.
+We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics.
+ADE refers to the average displacement error of the predicted trajectory $\widehat{pos}_{\tau+1:T}$ compared to the ground truth trajectory $pos_{\tau+1:T}$, and FDE represents the displacement error of the predicted goal $\widehat{pos}_{T}$ compared to the ground truth goal $pos_{T}$.
+$ADE=\frac{1}{T-\tau}\sum_{t=\tau+1}^{T}\left\|\widehat{pos}_{t}-pos_{t}\right\|$ (6)
+$FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (7)
+where $\tau$ represents the last time step of the observed trajectory; $T$ represents the last time step of the predicted trajectory; $\widehat{pos}_{t}$ and $pos_{t}$ represent the predicted coordinates and ground truth coordinates at time step $t$, respectively.
+Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE) metric, which will be discussed in detail in Section V-B.
+Implementation Details.
+In our model, the Transformer component adopts the original Transformer architecture [35].
+For the Neighbor Encoder module, the hidden size is set to 128.
+In the Seq-CVAE module, the LSTM hidden size and the latent size are configured to 256, and the MLP parameter design is illustrated in Figure 4.
+In the data preprocessing stage, the neighbor radius $r$ for social interaction is set to 2 meters, which is determined based on the average ADE performance.
+Due to significant variations in the walking speed and direction across different scenes, this study employs data augmentation by rotating trajectories at 0°, 90°, 180°, and 270°, along with scaling the trajectory length by a factor of 2.
+Notably, the direction-based interaction information $\mathcal{N}$ is transformed accordingly during rotation.
+To ensure the stability of the data distribution, all trajectories are shifted so that they start at coordinate (0,0).
+For the distance features in social interaction information, the reciprocal of the distance is taken, meaning the diminishing influence of pedestrians farther away.
+These values are subsequently normalized using Min-Max normalization.
+During training, the main components of InSyn (Interaction Encoder and Trajectory Generator) and the Seq-CVAE Goal Sampler are trained independently, and the teacher forcing technique is applied to facilitate training.
+For the Encoder-Generator, the SSOS strategy is applied.
+The learning rate is set to $1\times 10^{-4}$.
+For the Seq-CVAE, the KL divergence weight in the loss function is configured to 5 to balance the model’s reconstruction capability and latent space distribution requirements, and the learning rate is set to $1\times 10^{-3}$.
+The model is optimized using the Adam optimizer with 50 epochs.
+Mean Squared Error (MSE) serves as the loss function.
+During testing, the model generates predictions autoregressively.
+All experiments are conducted on a single NVIDIA RTX 4090 GPU, hosted on the AutoDL cloud platform.
+V-A Comparison
+V-A1 Comparison Baselines
 We compare our model with representative methods PECNet [7] and TF [11], as well as the latest work [42, 4, 5, 9].
 Additionally, we include some common baselines such as SoPhie [39] and CGNS [40].
-The baseline results are from officially-reported metrics.
+The baseline results are from officially reported metrics.
 Among these, SoPhie [39], CGNS [40], STG-DAT [27], and Goal-CurveNet [42] leverage visual information, which our approach does not utilize.
-TF [9] is the first to apply Transformers for human trajectory prediction; however, unlike our method, it focuses on individual motion without explicitly modeling social or environmental interactions.
-5.1.2 Evaluation
-The comparative results are demonstrated in Table  1.
+TF [11] is the first to apply Transformers for human trajectory prediction; however, unlike our method, it focuses on individual motion without considering social interactions.
+V-A2 Evaluation
+The comparative results are demonstrated in Table I.
 Compared to the aforementioned methods, our model achieves the best average ADE of 0.26.
@@
 Sparse Scenarios.
-However, on the Hotel dataset, our model underperforms relatively compared to original Transformer [9].
+However, on the Hotel dataset, our model underperforms relatively compared to the original Transformer [11].
 The Hotel dataset is a smaller scene area, which limits the effectiveness of InSyn’s distance-based interaction region partitioning.
@@
 In this scenario, the basic attention mechanism of the original Transformer is sufficient to capture motion patterns effectively, and the additional complex interaction modeling in our model may lead to overfitting, causing excessive reliance on interaction behaviors.
-This issue is also shared by other complex models such as [32] and [33], as shown in Table 1.
+This issue is also shared by other complex models such as [4] and [42], as shown in Table I.
 Overall, our model achieves the best average ADE among the compared models and performs particularly well in scenarios with more complex interactions.
 This demonstrates the model’s ability to effectively handle intricate interactions among pedestrians.
-5.2 Ablation Studies
-5.2.1 Interaction-Related Components Analysis
-For the interaction part, region partition divides the circular area centered on the agent’s position into four regions to capture directional interaction effects.
+V-B Ablation Studies
+V-B1 Interaction-Related Components Analysis
+For the interaction part, the region partition divides the circular area centered on the agent’s position into four regions to capture directional interaction effects.
 The interaction state, proposed in this study, categorizes interactions into three types: In Sync, Conflict, and No Neighbor.
-To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table 2.
+To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table II.
 These variations allow us to isolate the individual and combined effects of region partition and interaction state.
@@
 Case studies on the variations are provided in the supplementary material.
-From the results in Table  2, it is evident that the model incorporating both region rartition and interaction state, namely the original InSyn, achieves the best performance.
+From the results in Table  II, it is evident that the model incorporating both region partition and interaction state, namely the original InSyn, achieves the best performance.
 This demonstrates that explicitly dividing interaction behaviors enhances the model’s ability to extract complex interaction patterns among an agent’s neighbors, leading to more reasonable trajectory predictions.
 Specifically, in the absence of region partition, the model may struggle to discern the direction of influence, relying solely on training data distributions to infer future trajectory deviations—a limitation that undermines its generalizability.
-Similarly, if without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance.
+Similarly, without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance.
 This would underperform when the training and testing sets differ.
 For instance, if the training set contains more In Sync scenarios, the trained model may underestimate repulsive effects in Conflict situations.
-Table 2: ADE Result of Ablation Study on Region Partition and Interaction State.
-RP: region partition; IS: interaction state.
-w/o denotes Without.
-InSyn incorporates both region partition and interaction state, whereas the Baseline model includes neither.
+TABLE II: ADE Result of Ablation Study on Region Partition (RP) and Interaction State (IS)
 Methods ADE
@@
 Baseline 0.60 0.28 0.41 0.29 0.24 0.36
-5.2.2 SSOS Strategy Analysis
+V-B2 SSOS Strategy Analysis
 Additionally, we conducted an ablation study on the SSOS strategy proposed in this paper.
-SSOS is designed to mitigate the divergence at the first step when using Transformer’s Encoder-Decoder architecture for numerical time-series prediction tasks, such as pedestrian trajectory prediction.
+SSOS is designed to mitigate the divergence at the first step when using the Transformer’s Encoder-Decoder architecture for numerical time series prediction tasks, such as pedestrian trajectory prediction.
 In this context, steps $0:\tau$ represent observed input to the encoder, while time steps $\tau+1:T$ correspond to the future trajectory to be predicted.
@@
 where $\tau+1$ denotes the first prediction time step.
-Table  3 compares the average IDE results across all datasets using SSOS and SOS (set to $pos_{\tau}$).
+Table III shows the average ADE results across all datasets, comparing different ablations of our model.
+”w/o” refers to ”without” a component.
+The full model (InSyn) includes both RP and IS, while the Baseline model excludes both components.
 The results show that the SSOS strategy reduces the average IDE by approximately 6.58% compared to the SOS strategy.
@@
 Additionally, it mitigates error accumulation during autoregressive prediction.
-Table 3: IDE Result of Ablation Study on SSOS. InSyn-SSOS: trained using SSOS; InSyn-SOS: trained using SOS. InSyn-SSOS represents the original InSyn model.
+TABLE III: IDE Result of Ablation Study on SSOS
 Methods IDE
@@
 InSyn-SOS 0.118 0.051 0.085 0.062 0.065 0.076
-5.3 Visualization
-To further explore its practical performance, we conducted a qualitative analysis of the model’s behavior in real-world scenarios via visualization.
-Figure 6 illustrates the walking state of a specific agent (pedestrian 1) during the observation time steps, with details on its trajectory and interactions.
-After feeding this input into InSyn, the model outputs the predicted trajectory distribution, which is visualized using Kernel Density Estimation (KDE).
-In this example, pedestrian 2 moves in the same direction as the agent, suggesting synchronization.
-With the exception of a brief Conflict at $t=3$, when pedestrian 2 temporarily moves to the upper-right of the agent, pedestrian 2 remains largely In Sync with the agent.
-Consequently, despite consistently staying to the agent’s vicinity, pedestrian 2 exerts relatively weak influence on the agent’s trajectory.
-At $t=7$, pedestrian 3 enters the agent’s neighbor region while moving in the opposite direction.
-The model encodes this information—"a Conflict from the upper-left region at $t=7$, at a distance of $d_{t=7}^{lu}$"—into a representation.
-The LSTM aggregates representations across history $0:\tau$ ($\tau=7$) and the Trajectory Generator interprets its hidden state $h_{\tau}$ as an "avoidance" signal.
-As a result, during the initial prediction steps, the agent turns to the right to avoid pedestrian 3.
-Since the agent consistently walks toward the upper-right during the observation period, the goals sampled by Seq-CVAE are predominantly directed toward the upper-right.
-After predicting the "avoidance" behavior, the goal-driven attention mechanism guides the model to refocus on moving toward the sampled goal.
-Due to the inherent randomness in pedestrian trajectories (e.g., variations in speed), long-term goal predictions may occasionally deviate from the ground truth.
-However, intermediate trajectory predictions, including directional changes, remain relatively accurate (see Figure 6).
-Through visualizing pedestrian trajectories, the model demonstrates its sensitivity to interaction relationships and its ability to effectively leverage goal-driven predictions.
-More case studies are provided in the supplementary material.
-Figure 6: Visualization of Predicted Trajectory Distributions.
-We visualized the trajectory prediction of an agent (pedestrian 1) in a scene where In Sync, Conflict, and No Neighbor occur.
-The number of samples is $K$=20, meaning 20 goals were sampled.
-For each goal, a trajectory was generated in a goal-driven manner.
-Since pedestrian 4 remains outside the agent’s neighbor region, the model disregards its influence.
-6 Conclusion
-This paper introduces InSyn, a novel model that extends the Transformer architecture to explicitly capture complex pedestrian interaction patterns.
-Experimental results demonstrate that our model excels in scenarios involving intricate interaction behaviors, effectively modeling detailed temporal and spatial relationships.
-Furthermore, to mitigate initial prediction divergence when applying the Transformer’s encoder-decoder framework to numerical sequence data, we propose the SSOS strategy, which relatively alleviates this challenge.
+V-C Case Study
+Figure 5: Case Study of Region Partition and Interaction State.
+This figure compares three variants of InSyn: (1) without region partition (w/o-RP), (2) without interaction state (w/o-IS), and (3) the full model incorporating both components (InSyn).
+To isolate the effect of interaction modeling, we exclude the Goal Sampler in this evaluation.
+To further investigate the effectiveness of region partition and interaction state, we conduct a Case study.
+The cases (see Figure 5) cover various interaction scenarios, including paired walking, group walking, and pedestrian conflict.
+In Sync Scenarios.
+Cases (a), (b), (e), and (f) represent instances where pedestrians walk in coordination with others nearby.
+In cases (a) and (b), where the target pedestrian moves alongside a companion, InSyn preserves the direction and velocity of the agent’s motion at the end of the observation period, closely matching the ground truth.
+In contrast, both ablated variants produce noticeable deviations.
+These cases suggest that the influence of the neighboring pedestrian is relatively weak.
+InSyn effectively models the behaviors by accurately capturing the reduced impact of the neighbor on the agent’s motion.
+Case (e) depicts a group walking situation with multiple neighboring pedestrians moving in sync.
+Here, only InSyn predicts the correct direction of motion, while both variants mistakenly infer an opposite trajectory, highlighting their failure to capture the low-impact nature of synchronized group behavior.
+Case (f) presents a simple, slow-moving scenario where all models achieve comparable performance, showing that the base model alone suffices for simple prediction tasks.
+Conflict Scenarios.
+Cases (c) and (d) illustrate interactions where conflict plays a dominant role.
+In case (c), although the agent’s last observed motion trends slightly upward, InSyn predicts a flatter, rightward trajectory that better aligns with the ground truth.
+This adjustment reflects the model’s ability to recognize a conflicting pedestrian approaching from above, thereby reducing unnecessary upward movement.
+The w/o-RP and w/o-IS variants, lacking directional awareness and interaction-type recognition respectively, simply extrapolate the upward motion and thus produce larger errors.
+Case (d), however, reveals a limitation of InSyn.
+In this instance, the model appears to overestimate the influence of a pedestrian approaching from below, leading to an excessive upward deviation from the true path.
+This suggests that while InSyn improves reliability and interpretability in most complex scenarios, further refinement is needed to enhance control precision in handling conflict interactions.
+These case studies demonstrate that modeling explicit interaction types and directional influence contributes to more trustworthy and socially-aware predictions.
+Compared to black-box representations, our approach enhances interpretability, allowing for more transparent reasoning about how different pedestrian behaviors shape trajectory outcomes.
+This interpretability not only builds trust in the model’s decisions but also provides actionable insights for refining interaction modeling strategies.
+VI Conclusion
+This paper presents InSyn, a trustworthy trajectory prediction model that extends the Transformer architecture to explicitly capture pedestrian complex interaction patterns.
+By introducing interpretable structure into otherwise black-box models, InSyn improves interpretability and robustness in socially complex environments.
+Experimental results validate the effectiveness of our approach in both dense and sparse scenarios, offering more reliable predictions.
+Besides, our proposed SSOS strategy enhances prediction accuracy by mitigating initial-step divergence in numerical sequence modeling.
+A detailed case study further illustrates how modeling explicit interaction patterns contributes to interpretable and trustworthy predictions in diverse real-world scenarios.
 Future work will focus on two key directions.
 First, we aim to design a more refined approach for partitioning interaction regions to better handle challenges posed by small-scale scenes.
-Second, we plan to apply SSOS strategy to other types of numerical sequential data beyond trajectory prediction, exploring its potential in diverse domains.
-{ack}
+Second, we plan to apply the SSOS strategy to other types of numerical sequential data beyond trajectory prediction, exploring its potential in diverse domains.
+Acknowledgment
+This work was supported by the Natural Science Foundation of Shanghai (No. 23ZR1422800). The authors would like to thank the SusCom Lab and Dr. Taowen Wang at HKUST(GZ) for their valuable support.

### v3

@@
 Abstract
-Accurate and reliable pedestrian trajectory prediction is critical for the safety and robustness of intelligent applications, yet achieving trustworthy prediction remains highly challenging due to the complexity of interactions among pedestrians.
-Previous methods often adopt black-box modeling of pedestrian interactions, treating all neighbors uniformly.
-Despite their strong performance, such opaque modeling limits the reliability of predictions in safety-critical real-world deployments.
+Accurate and reliable pedestrian trajectory prediction is critical for the application of intelligent applications, yet achieving trustworthy prediction remains highly challenging due to the complexity of interactions among pedestrians.
+Previous methods often adopt black-box modeling of pedestrian interactions.
+Despite their strong performance, such opaque modeling limits the reliability of predictions in real-world deployments.
 To address this issue, we propose InSyn (Interaction-Synchronization Network), a novel Transformer-based model that explicitly captures diverse interaction patterns (e.g., walking in sync or conflicting) while effectively modeling direction-sensitive social behaviors.
 Additionally, we introduce a training strategy, termed Seq-Start of Seq (SSOS), designed to alleviate the common issue of initial-step divergence in numerical time-series prediction.
-Experiments on the ETH and UCY datasets demonstrate that our model not only outperforms recent black-box baselines in prediction accuracy, especially under high-density scenarios, but also provides stronger interpretability, achieving a favorable trade-off between reliability and accuracy.
+Experiments on the ETH and UCY datasets demonstrate that our model not only outperforms recent black-box baselines in prediction accuracy, especially under high-density scenarios, but also provides transparent interaction modeling, as shown in the case study.
 Furthermore, the SSOS strategy proves to be effective in improving sequential prediction performance, reducing the initial-step prediction error by approximately 6.58%.
-Index Terms:
-Human Trajectory Prediction, Interpretable Modeling, Attention Mechanism
-I Introduction
-Pedestrian trajectory prediction is essential for safety-critical applications such as autonomous driving [1] and robotic navigation [2].
-Accurate and reliable forecasting enables intelligent systems to understand human behavior better and ensures safe and trustworthy predictions.
+Code is avaliable at https://github.com/rickzky1001/InSyn
+1 Introduction
+Pedestrian trajectory prediction is essential for real-world applications such as autonomous driving [18] and robotic navigation [3].
+Accurate and reliable forecasting enables intelligent systems to understand human behavior better and ensures trustworthy predictions.
 However, the presence of complex pedestrian interactions poses significant challenges to the task.
-In a given scene, pedestrians may respond to nearby individuals in different ways.
+In a given scene, pedestrians may respond to nearby individuals in different ways [13].
 For instance, the trajectories of two pedestrians may exhibit conflict.
@@
 F represents the interaction effect between pedestrians.
-In traditional approaches (top), all neighbors of the agent are treated as being in the same state, with interactions solely based on the distance between the agent and its neighbors.
-This simplistic assumption may lead to overestimating the influence of fellow pedestrians.
-Our method (bottom) introduces a more refined modeling strategy by considering the specific states of neighboring pedestrians (e.g., In Sync or Conflict), providing a more nuanced understanding of the neighboring interaction.
-Recently, extensive studies have investigated pedestrian interactions [3, 4, 5, 6], using approaches like social pooling layers [3, 7], social force mechanism [6], and graph neural networks (GNNs) [8, 9].
-However, a common characteristic of these methods is that they rely on relative positions between an agent and its neighbors to model influence, treat all neighboring pedestrians by black-box representation, failing to differentiate patterns such as In Sync and Conflict (see Figure 1).
+In traditional approaches (top), all neighbors of the agent are treated as being in the same state.
+Our method (bottom) introduces a more refined modeling strategy by considering the specific states of neighboring pedestrians, providing a more nuanced understanding of the neighboring interaction.
+Recently, extensive studies have investigated pedestrian interactions [1, 31, 23, 19, 40], using approaches like social pooling layers [1, 6], social force mechanism [40, 41], and graph neural networks (GNNs) [38, 17, 4].
+However, a common characteristic of these methods is that they rely on features like relative position between an agent and its neighbors to model influence, treat all neighboring pedestrians by black-box representation, failing to differentiate patterns such as In Sync and Conflict (see Figure 1).
 This oversimplification may lead to overfitting and untrustworthy predictions.
-For example, ignoring synchronized walking (e.g., friends walking side-by-side) can lead to overlapping predicted paths.
-In conflict cases, such as head-on encounters, the absence of explicit modeling may cause unrealistic predictions that either overestimate collisions or miss necessary evasive actions.
-This black-box modeling not only reduces prediction accuracy, but also limits interpretability and trustworthiness in safety-critical scenarios such as machine navigation and intelligent transportation systems.
+For example, failing to recognize synchronized walking (e.g., friends walking side-by-side) and mistakenly modeling them with repulsive forces can lead to unrealistic divergence in their predicted paths.
+This black-box modeling not only reduces prediction accuracy, but also limits interpretability and trustworthiness in real-world scenarios.
 These limitations become particularly pronounced in crowded or dynamic environments.
-To address this issue, we need a method that understands what’s exactly going on between pedestrians—recognizing concrete interaction patterns rather than treating them as abstract black-box inputs—to ensure more reliable and interpretable predictions (see Figure 1).
+To address this issue, we need a method that understands what’s exactly going on between pedestrians—recognizing interaction patterns rather than treating them as abstract black-box inputs—to ensure more reliable and accurate predictions (see Figure 1).
 Although pedestrian trajectories are highly stochastic, they usually have a clear goal [27].
-By focusing on the target goal, the performance of the model can be significantly enhanced.
-Experimental results from Giuliari et al.
-[11] demonstrate that the attention mechanism is well-suited for processing sequential data with irregular time intervals (e.g., observed trajectories and goals), effectively capturing long-term dependencies in pedestrian motion and providing reliable goal-driven functionality.
-This paper proposes the InSyn, a Transformer-based model.
+Focusing on the target goal can significantly enhance model performance.
+Building on this insight and the suitability of the attention mechanism for capturing long-range dependencies in sequential data [7], we propose InSyn, a Transformer-based model.
 The model consists of three components: (1) Interaction Encoder: designed to explicitly extract interaction information and integrates interaction features, goal-driven behavior, and observed trajectories through the self-attention mechanism.
-(2) Trajectory Generator: forms an attention mechanism in conjunction with the Interaction Encoder, and incorporates the proposed SSOS strategy to alleviate divergence at the initial prediction step.
+(2) Trajectory Generator: forms an cross-attention mechanism in conjunction with the Interaction Encoder, and incorporates the proposed SSOS strategy to alleviate divergence at the initial prediction step.
 (3) Seq-CVAE Goal Sampler: a conditional generative model specifically designed for sequential prediction, used for goal sampling.
-We evaluate our model on the well-established pedestrian trajectory prediction datasets ETH [12] and UCY [13].
-The experimental results demonstrate that our model outperforms recent methods in key metrics, particularly in scenarios with complex interactions.
-We further conduct ablation studies to show the superiority of interaction modeling and the SSOS strategy.
-Additionally, the case study highlights the interpretability of our approach by illustrating how specific interaction patterns influence prediction outcomes.
-Figure 2: Left is the illustration of the input $S_{0:\tau}$.
-At each time step, the walking state $S_{k}$ comprises the 2D coordinates $(x,y)$ and interaction information $N_{k}$.
-Right demonstrates the scenarios of In Sync, Conflict and No Neighbor state.
 In summary, our work makes two key contributions:
@@
 This strategy, termed SSOS, mitigates the divergence in the first prediction step, thereby reducing error accumulation and improving overall performance.
-II Related Work
-II-A Pedestrian Trajectory Prediction
-Pedestrian trajectory prediction methods can be categorized into traditional machine learning approaches and deep learning-based approaches.
-Traditional methods, such as Markov decision processes [14] and Gaussian distributions [15], are efficient in handling simple scenarios.
-However, they often struggle with complex data distributions and exhibit limited generalization capabilities.
-In contrast, deep learning approaches have shown significant improvements in modeling complex patterns [16, 17, 18, 19].
-Long Short-Term Memory (LSTM) network [20], a classic time-series model, is applied in this field.
-Social-LSTM [3] introduced a pooling mechanism to incorporate interactions into trajectory prediction.
-Yang et al.
-[21] proposed a variant LSTM structure for information sharing within scenes, while SR-LSTM [22] combined neighbor intentions and message passing to enhance prediction accuracy.
-These works leverage the strong temporal feature extraction capabilities of LSTM.
-In recent years, Transformers have gained popularity in trajectory prediction due to their ability to capture long-range dependencies.
-Liu et al.
-[23] combined CNN and Transformer, using the attention mechanism to process interaction information.
-MRGTraj [24] integrated a Transformer encoder with a novel decoder, achieving trajectory prediction through a mapping-refining-generating structure.
-CITraNet [25] significantly improved prediction accuracy by introducing an innovative Transformer-based Gumbel distribution network.
-Building on these advancements, our research combines the strengths of LSTM and Transformer.
-By leveraging the temporal feature extraction capabilities of LSTM, we enhance the Transformer’s attention mechanism with processed temporal interaction features, thereby improving prediction performance.
-II-B Pedestrian Interaction Modeling
+2 Related Work
 Pedestrian interaction is a key factor influencing prediction accuracy. Pedestrian motion is not only driven by individual goals but also affected by interactions with surrounding pedestrians.
-Current mainstream methods for interaction modeling include social pooling layers [3, 26], GNNs [9, 8], and attention mechanisms [27, 28].
+Current mainstream methods for interaction modeling include social pooling layers [1, 34, 10], GNNs [17, 38, 24, 29], and attention mechanisms [16, 35, 20].
 Among these, the social attention module aggregates interaction information by analyzing correlations between pedestrian motion and future trajectories [34].
@@
 However, most of these methods rely on black-box representations of social influence.
-This limits the model’s ability to produce reliable and interpretable predictions, especially in complex or safety-critical scenarios.
-In contrast, our approach explicitly captures structured interaction patterns, such as walking In Sync or engaging in Conflict, aiming to enhance the trustworthiness and transparency of socially-aware trajectory prediction.
-III Problem Formulation
+This limits the model’s ability to produce reliable and interpretable predictions.
+In contrast, our approach explicitly leverages structured interaction patterns, such as walking In Sync or engaging in Conflict, aiming to enhance the trustworthiness and transparency of socially-aware trajectory prediction.
+3 Problem Formulation
+Figure 2: Left is the illustration of the input $S_{0:\tau}$.
+At each time step, the walking state $S_{k}$ comprises the 2D coordinates $(x,y)$ and interaction information $N_{k}$.
+Right demonstrates the scenarios of In Sync, Conflict and No Neighbor state.
+Note that the 4-regions partition and the scenarios identification use simple spatial-temporal rules for transparency, more complex methods could be applied here but beyond our scope.
 In this study, pedestrian trajectory prediction is formulated as follows: given the walking states $S_{0:\tau}$ of a certain agent during the observation time steps $\mathord{0:\tau}$, the model outputs the predicted trajectory $\widehat{pos}_{\tau+1:T}$ of the agent for the future time steps $\mathord{\tau+1:T}$.
@@
 Each walking state $S_{t}^{p}$ consists of the agent’s position $pos_{t}=(x_{t},y_{t})$ in 2D space and the pedestrian interaction information $N_{t}$.
-Since interactions are directional, we adopt a region partition strategy around each agent.
+Since interactions are directional, we adopt a region partition strategy [1] around each agent.
 The interaction information $N_{t}$ is defined as
-$N_{t}=\left[\left(f_{t}^{lu},d_{t}^{lu}\right),\left(f_{t}^{ru},d_{t}^{ru}\right),\left(f_{t}^{ld},d_{t}^{ld}\right),\left(f_{t}^{rd},d_{t}^{rd}\right)\right]$ (1)
-where $f_{t}^{lu},f_{t}^{ru},f_{t}^{ld},f_{t}^{rd}$ represent the interaction states in the left-up, right-up, left-down, and right-down regions, respectively.
-These patterns include No Neighbor, In Sync, and Conflict, as shown in Figure 2.
-The construction of interaction patterns is highly transparent and trustworthy, as it relies on interpretable spatial-temporal neighbor changes.
+$N_{t}=\left\{(f_{t}^{r},d_{t}^{r})\right\}_{r\in\{\mathrm{lu,ru,ld,rd}\}}$ (1)
+where $f_{t}^{r}$ represents the interaction state in region $r$ at time $t$; $d_{t}^{r}$ indicates the distances to the nearest neighbor; $\{\mathrm{lu,ru,ld,rd}\}$ represent the left-up, right-up, left-down, right-down of the agent respectively.
+The interaction state $f_{t}^{r}$ can be No Neighbor, In Sync, and Conflict, as shown in Figure 2.
+The construction of interaction patterns is highly transparent.
 Specifically, if the nearest neighbor in a given region at time $t$ is the same as the one at time t1t-1, the interaction state is classified as In Sync; otherwise, it is marked as Conflict.
-In cases of No Neighbor or In Sync, the interaction influence of the neighbor on the agent’s future trajectory may be negligible.
+A region with no pedestrians is classified as No Neighbor, and a large distance value is assigned to indicate minimal influence.
+In cases of No Neighbor or In Sync, the interaction influence of the neighbor on the agent’s future trajectory may be limited.
 However, if a new pedestrian suddenly enters a region, referred to as Conflict, it is more likely to impact the agent’s trajectory.
-The terms $d_{t}^{lu},d_{t}^{ru},d_{t}^{ld},d_{t}^{rd}$ indicate the distances to the nearest neighbor in each region.
-A region with no pedestrians is classified as No Neighbor, and a large distance value is assigned to indicate minimal influence.
-The goal of this study is to train a generative model
-$m_{\theta}\left(\widehat{pos}_{\tau+1:T}\mid S_{0:\tau}\right)$ (2)
-where $\theta$ represents the model parameters.
-The model aims to predict the distribution of pedestrians’ future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
+This identification uses simple rules for transparency and it could be replaced by a model-based method.
+The goal of this study is to train a generative model $m_{\theta}\left(\widehat{pos}_{\tau+1:T}\mid S_{0:\tau}\right)$, aiming to predict the distribution of pedestrians’ future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
+4 Methodology
 Figure 3: Overview of the InSyn framework for trajectory prediction.
@@
 The input observed walking state includes the agent’s trajectory positions $pos_{0:\tau}$ and its interaction information $N_{0:\tau}$ within the observed time $0:\tau$.
-IV Methodology
 In this section, we propose InSyn, a Transformer-based model tailored for the interpretable modeling of specific social interactions.
@@
 The framework of InSyn is illustrated in Figure 3, and the Seq-CVAE is depicted in Figure 4.
-We first discuss how our model captures specific interactions and achieves goal-driven functionality.
-We then elaborate on the motivation behind the SSOS strategy and, finally, the design of the Seq-CVAE module.
-IV-A Interaction Encoder
+4.1 Interaction Encoder
 The Interaction Encoder integrates observed trajectory, interaction information, and sampled goal to extract spatiotemporal features.
 These features are utilized for goal-driven and interaction-sensitive trajectory prediction.
-IV-A1 Goal-Driven
-Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  IV-B) is combined with the position information through concatenation.
+4.1.1 Goal-Driven.
+Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  4.3) is combined with the position information through concatenation.
 Notably, since self-attention cannot inherently capture positional information, positional encoding is used to encode the input positions.
 This encoding is based on the sequence positions of the inputs.
-Therefore, before concatenating $pos_{0:\tau}^{p}$ with the sampled goal $\widehat{pos}_{T}$, padding needs to be inserted between them to ensure that the goal is assigned the corresponding positional encoding.
-IV-A2 Interaction
-To facilitate trustworthy modeling of social interactions, we introduce a modular neighbor encoder that encodes pattern-aware and distance-based interaction effects.
+Therefore, to ensure the sampled goal $\widehat{pos}_{T}$ is assigned the correct positional encoding, a $T-\tau-1$ length padding tokens must be inserted between the observed trajectory $pos{0:\tau}^{p}$ and the goal before concatenation.
+4.1.2 Interaction.
+To facilitate trustworthy modeling of social interactions, we introduce a modular neighbor encoder that encodes pattern-aware effects.
 The interaction patterns between pedestrians are extracted through a neighbor encoder module followed by an LSTM.
-The agent $p$’s social interaction information, represented as $\mathcal{N}^{p}=\left(N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}\right)$, are encoded using the neighbor encoder.
-which consists of three components: an embedding layer, a Distance Gate, and a Social Pooling Layer.
-The embedding layer maps interaction states (No Neighbor, In Sync, Conflict) to learnable representations.
+The agent $p$’s observed social interaction information, represented as $\mathcal{N}^{p}=\left(N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}\right)$, are encoded using the neighbor encoder.
+which consists of three components: an embedding layer, a Distance Gate, and a Social Pooling Layer, as shown in Figure 3.
+The embedding layer maps interaction states $f_{t}^{r}$ to learnable representations.
 These representations are then combined with the processed distance $d$ through element-wise multiplication.
@@
 Consequently, $h_{\tau}$ is directly concatenated with the positionally encoded representation (see Figure 3).
-IV-B SSOS: Seq-Start of Seq
-[…]
-Nevertheless, we note that when the Transformer’s encoder-decoder architecture is applied to trajectory prediction—a numerical time-series prediction task—using a single value as the SOS can lead to a lack of smoothness in the transition between the first predicted value $\widehat{pos}_{\tau+1}$ and the observed trajectory $pos_{0:\tau}$.
-That is to say, $\widehat{pos}_{\tau+1}$ may deviate noticeably from the ground truth $pos_{\tau+1}$.
+4.2 Trajectory Generator
+The Trajectory Generator incorporates the SSOS Strategy into the Transformer Decoder [28] to alleviate initial prediction divergence.
+The Transformer’s decoder relies on the Start of Sequence (SOS) as the initial input when generating the first output.
+In Natural Language Processing (NLP), the Begin-of-Sentence token $\left<bos\right$> is commonly used as the SOS, which aligns with the semantic structure of text.
+However, in sequential trajectory prediction, the selection of SOS should consider the spatial information; otherwise, it may introduce noise.
+Giuliari et al.
+[7] set the SOS to $(0,0)$ for trajectory prediction.
+While $(0,0)$ represents a fixed position in spatial coordinates, using it as the SOS may introduce noise and mislead the model due to its lack of alignment with the observed trajectory.
+Therefore, it is reasonable to consider using the position at the last observed time step, $\widehat{pos}_{\tau}$, as the SOS token of the decoder.
+Nevertheless, we note that when the Transformer’s encoder-decoder architecture is applied to trajectory prediction, a numerical time-series prediction task, using a single value as the SOS can lead to a lack of smoothness in the transition between the first predicted value $\widehat{pos}_{\tau+1}$ and the observed trajectory $pos_{0:\tau}$.
+That is to say, $\widehat{pos}_{\tau+1}$ may deviate noticeably from the last observation point $pos_{\tau}$.
 This issue likely arises because, during the initial decoding stage, the Transformer decoder relies solely on the self-attention with the SOS and the cross-attention with the encoder’s output.
 Insufficient information or excessive noise in the SOS may lead to alignment bias, thus reducing prediction accuracy.
-To mitigate this issue, we propose using the sequence data $pos_{0,1,...,\tau}$ as the SOS.
+To mitigate this issue, we propose using the sequence data $pos_{0:\tau}$ as the SOS.
 We refer to this approach as SSOS and use it during training.
@@
 $\quad+\lambda_{2}\cdot\text{MSE}(\widehat{pos}_{\tau+1:T-1},pos_{\tau+1:T-1})$ (2)
-where $\widehat{pos}_{1:\tau}$ denotes the reconstructed observed trajectory from time $1$ to $\tau$; $\widehat{pos}_{\tau+1:T-1}$ denotes the predicted trajectory excluding the sampled goal; $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that balance the contributions of the reconstruction loss and the prediction loss; MSE is the Mean Squared Error.
-However, the SSOS strategy generates additional outputs $\widehat{pos}_{1:\tau}$, which are not part of the target prediction. These byproducts may introduce additional computational overhead.
-IV-C Seq-CVAE: LSTM-Based CVAE
-RNN (typically LSTM [20]) and Conditional Variational Autoencoder (CVAE) [36] have been widely applied in pedestrian trajectory prediction [37, 21, 6, 3].
-We combine the two approaches and propose a generator, Seq-CVAE (Sequence CVAE), which is designed for sampling trajectory goals.
+where $\widehat{pos}_{1:\tau}$ denotes the reconstructed observed trajectory from time $1$ to $\tau$; $\widehat{pos}_{\tau+1:T-1}$ denotes the predicted trajectory excluding the sampled goal; $\lambda_{1}$ and $\lambda_{2}$ are used for balancing.
+Ablation results in Section 5.2 shows that this strategy enchances the performance of the model.
+However, the SSOS strategy generates additional outputs $\widehat{pos}_{1:\tau}$, which are not part of the target prediction.
+These byproducts may introduce additional computational overhead.
+4.3 Seq-CVAE Goal Sampler
+RNN (typically LSTM [9]) and Conditional Variational Autoencoder (CVAE) [25, 12] have been widely applied in pedestrian trajectory prediction [37, 40, 1].
+We combine the two approaches and propose a sequence-based generator, Seq-CVAE (Sequence CVAE), which is designed for sampling trajectory goals.
 The model predicts goals given the condition $c$, which is derived from the observed trajectory $pos_{0:\tau}$.
@@
 Therefore, for the Seq-CVAE, the condition $c$ for the encoder and decoder is defined as
-$c_{Encoder}$ $=\mathrm{\textit{FC}}(\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\}))$ (4)
-$c_{Decoder}$ $=\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\})$ (5)
+$c_{Encoder}=\mathrm{\textit{FC}}(\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\})),\quad c_{Decoder}=\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\})$ (3)
 where $FC$ represents the fully connected layer, $h_{\tau}$ denotes the hidden state of LSTM at the last observed time step $\tau$, and $pos_{0:\tau}$ represents the coordinates of the observed trajectory.
 […]
-V Experiment
-TABLE I: Results comparison on all datasets. Bold: Best
+5 Experiment
+5.0.1 Datasets and Evaluation Protocol.
+We evaluate on the standard ETH [21] and UCY [14] datasets, following the leave-one-out strategy [30, 40, 39, 5].
+As in prior work [37, 38, 39], we do not use visual information.
+The model observes 3.2 seconds (8 frames at 2.5 fps) and predicts the next 4.8 seconds.
+Evaluation employs the standard best-of-$K=20$ protocol [15, 40, 6, 38] for fair comparison.
+5.0.2 Metrics.
+We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics.
+$ADE=\frac{1}{T-\tau}\sum_{t=\tau+1}^{T}\left\|\widehat{pos}_{t}-pos_{t}\right\|,\quad FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (4)
+where $\tau$ represents the last time step of the observed trajectory; $T$ represents the last time step of the predicted trajectory;
+Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE), which will be discussed in Section 5.2.
+5.0.3 Implementation Details.
+Our model builds on the Transformer architecture [28].
+Key configurations: Neighbor Encoder hidden size 128; Seq-CVAE LSTM and latent sizes 256 (see Fig.
+4); neighbor radius 2m.
+We apply data augmentation (rotation and scaling) and normalize trajectories to start at (0,0).
+The Interaction Encoder/Generator ($lr=1\times 10^{-4}$) and Seq-CVAE ($lr=1\times 10^{-3}$, KL weight=5) are trained separately for 500 epochs with Adam and MSE loss.
+Testing process is autoregressive.
+We performed comprehensive data processing and augmentation; the detailed pipeline is provided in our code.
+5.1 Comparison
+5.1.1 Comparison Baselines.
+We compare our model with representative methods PECNet [6] and TF [7], as well as the rencent work [32, 31, 23, 17].
+Additionally, we include some common baselines such as SoPhie [22] and CGNS [15].
+The baseline results are from officially reported metrics.
+Among these, SoPhie [22], CGNS [15], STG-DAT [16], and Goal-CurveNet [32] leverage visual information, which our approach does not utilize.
+TF [7] is the first to apply Transformers for human trajectory prediction; however, unlike our method, it focuses on individual motion without considering social interactions.
+5.1.2 Evaluation.
+Table 1: Results comparison on all datasets. Bold: Best
 Methods Year ETH Hotel Univ Zara01 Zara02 Average
@@
 InSyn (Ours) - 0.36 0.77 0.27 0.47 0.31 0.54 0.20 0.44 0.15 0.36 0.26 0.52
-Datasets.
-We use the well-established datasets in the field of pedestrian trajectory prediction: ETH [12] and UCY [13].
-The datasets include five sub-datasets, covering four different scenarios and 1536 trajectories.
-These trajectories include various interaction scenarios such as walking together, as couples, and in groups.
-Similar to previous work [43, 6, 34], we adopt the leave-one-out strategy for splitting the training and testing sets.
-That is, four datasets are used for training and validation, and one dataset is used for testing.
-Since partitioning the neighboring regions requires real-world distances, we convert the coordinate scale to real-world metrics in meters.
-Due to the different frame rates of the ETH and UCY datasets, we unify both to 2.5 fps, corresponding to a time step of 0.4 seconds.
-Evaluation Protocol.
-Like many previous work [21, 8, 34], visual information is not used in our approach.
-During the test phase, our model takes the trajectory within 3.2 seconds as input and outputs the trajectory for the subsequent 4.8 seconds.
-For fair comparison, we adopt the best-of-$K$ protocol, which is sampling $K=20$ times and choosing the best result for metric computation.
-This is the standard protocol in pedestrian trajectory prediction [40, 6, 7, 8].
-Metrics.
-We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics.
-ADE refers to the average displacement error of the predicted trajectory $\widehat{pos}_{\tau+1:T}$ compared to the ground truth trajectory $pos_{\tau+1:T}$, and FDE represents the displacement error of the predicted goal $\widehat{pos}_{T}$ compared to the ground truth goal $pos_{T}$.
-$ADE=\frac{1}{T-\tau}\sum_{t=\tau+1}^{T}\left\|\widehat{pos}_{t}-pos_{t}\right\|$ (6)
-$FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (7)
-where $\tau$ represents the last time step of the observed trajectory; $T$ represents the last time step of the predicted trajectory; $\widehat{pos}_{t}$ and $pos_{t}$ represent the predicted coordinates and ground truth coordinates at time step $t$, respectively.
-Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE) metric, which will be discussed in detail in Section V-B.
-Implementation Details.
-In our model, the Transformer component adopts the original Transformer architecture [35].
-For the Neighbor Encoder module, the hidden size is set to 128.
-In the Seq-CVAE module, the LSTM hidden size and the latent size are configured to 256, and the MLP parameter design is illustrated in Figure 4.
-In the data preprocessing stage, the neighbor radius $r$ for social interaction is set to 2 meters, which is determined based on the average ADE performance.
-Due to significant variations in the walking speed and direction across different scenes, this study employs data augmentation by rotating trajectories at 0°, 90°, 180°, and 270°, along with scaling the trajectory length by a factor of 2.
-Notably, the direction-based interaction information $\mathcal{N}$ is transformed accordingly during rotation.
-To ensure the stability of the data distribution, all trajectories are shifted so that they start at coordinate (0,0).
-For the distance features in social interaction information, the reciprocal of the distance is taken, meaning the diminishing influence of pedestrians farther away.
-These values are subsequently normalized using Min-Max normalization.
-During training, the main components of InSyn (Interaction Encoder and Trajectory Generator) and the Seq-CVAE Goal Sampler are trained independently, and the teacher forcing technique is applied to facilitate training.
-For the Encoder-Generator, the SSOS strategy is applied.
-The learning rate is set to $1\times 10^{-4}$.
-For the Seq-CVAE, the KL divergence weight in the loss function is configured to 5 to balance the model’s reconstruction capability and latent space distribution requirements, and the learning rate is set to $1\times 10^{-3}$.
-The model is optimized using the Adam optimizer with 50 epochs.
-Mean Squared Error (MSE) serves as the loss function.
-During testing, the model generates predictions autoregressively.
-All experiments are conducted on a single NVIDIA RTX 4090 GPU, hosted on the AutoDL cloud platform.
-V-A Comparison
-V-A1 Comparison Baselines
-We compare our model with representative methods PECNet [7] and TF [11], as well as the latest work [42, 4, 5, 9].
-Additionally, we include some common baselines such as SoPhie [39] and CGNS [40].
-The baseline results are from officially reported metrics.
-Among these, SoPhie [39], CGNS [40], STG-DAT [27], and Goal-CurveNet [42] leverage visual information, which our approach does not utilize.
-TF [11] is the first to apply Transformers for human trajectory prediction; however, unlike our method, it focuses on individual motion without considering social interactions.
-V-A2 Evaluation
-The comparative results are demonstrated in Table I.
+The comparative results are demonstrated in Table 1.
 Compared to the aforementioned methods, our model achieves the best average ADE of 0.26.
-For goal prediction, the Seq-CVAE achieves excellent performance with an average FDE of 0.52, slightly trailing only behind PECNet [7] and TDAGCN [4] despite its relatively lightweight design.
-In contrast to FDE, which only evaluates the final predicted position, ADE measures error across all predicted trajectory points, with greater emphasis on intermediate trajectory deviations, thereby better reflecting the model’s capacity to model complex pedestrian interactions.
-To focus on the model’s ability to capture interactions, we adopt ADE as the evaluation metric for assessing and comparing performance across different scenarios.
+For goal prediction, the Seq-CVAE achieves an average FDE of 0.52, which is competitive though slightly higher than some state-of-the-art methods [6, 31].
+This is likely due to its separated design from the main interaction parts.
+Since ADE measures overall trajectory consistency and better reflects the model’s capacity to capture complex pedestrian interactions, we emphasize it as the primary metric for evaluation of interaction modeling.
 Crowded Scenarios.
@@
 In this scenario, the basic attention mechanism of the original Transformer is sufficient to capture motion patterns effectively, and the additional complex interaction modeling in our model may lead to overfitting, causing excessive reliance on interaction behaviors.
-This issue is also shared by other complex models such as [4] and [42], as shown in Table I.
+This issue is also shared by other complex models such as [31] and [32], as shown in Table 1.
 Overall, our model achieves the best average ADE among the compared models and performs particularly well in scenarios with more complex interactions.
 This demonstrates the model’s ability to effectively handle intricate interactions among pedestrians.
-V-B Ablation Studies
-V-B1 Interaction-Related Components Analysis
-For the interaction part, the region partition divides the circular area centered on the agent’s position into four regions to capture directional interaction effects.
+5.2 Ablation Study
+5.2.1 Interaction-Related Components Analysis.
+For the interaction part, the region partition divides the area centered on the agent into four regions to capture directional interaction effects.
 The interaction state, proposed in this study, categorizes interactions into three types: In Sync, Conflict, and No Neighbor.
-To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table II.
+To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table 2.
+“w/o” refers to “without” a component.
+The full model (InSyn) includes both RP and IS, while the Baseline model excludes both components.
 These variations allow us to isolate the individual and combined effects of region partition and interaction state.
-The original InSyn incorporates both region partition and interaction state.
-Case studies on the variations are provided in the supplementary material.
-From the results in Table  II, it is evident that the model incorporating both region partition and interaction state, namely the original InSyn, achieves the best performance.
-This demonstrates that explicitly dividing interaction behaviors enhances the model’s ability to extract complex interaction patterns among an agent’s neighbors, leading to more reasonable trajectory predictions.
-Specifically, in the absence of region partition, the model may struggle to discern the direction of influence, relying solely on training data distributions to infer future trajectory deviations—a limitation that undermines its generalizability.
-Similarly, without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance.
-This would underperform when the training and testing sets differ.
-For instance, if the training set contains more In Sync scenarios, the trained model may underestimate repulsive effects in Conflict situations.
-TABLE II: ADE Result of Ablation Study on Region Partition (RP) and Interaction State (IS)
+Table 2: ADE Result of Ablation Study on Region Partition (RP) and Interaction State (IS)
 Methods ADE
@@
 Baseline 0.60 0.28 0.41 0.29 0.24 0.36
-V-B2 SSOS Strategy Analysis
+It is evident that the model incorporating both region partition and interaction state, namely the original InSyn, achieves the best performance.
+This demonstrates that explicitly dividing interaction behaviors enhances the model’s ability to extract complex interaction patterns among an agent’s neighbors, leading to more reasonable trajectory predictions.
+Specifically, in the absence of region partition, the model may struggle to discern the direction of influence, relying solely on training data distributions to infer future trajectory deviations—a limitation that undermines its generalizability.
+Similarly, without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance and direction.
+This would underperform when the training and testing sets differ.
+For instance, if the training set contains more In Sync scenarios, the trained model may underestimate repulsive effects in Conflict situations.
+More case studies on the variations are provided in Section 5.3.
+5.2.2 SSOS Strategy Analysis.
 Additionally, we conducted an ablation study on the SSOS strategy proposed in this paper.
-SSOS is designed to mitigate the divergence at the first step when using the Transformer’s Encoder-Decoder architecture for numerical time series prediction tasks, such as pedestrian trajectory prediction.
-In this context, steps $0:\tau$ represent observed input to the encoder, while time steps $\tau+1:T$ correspond to the future trajectory to be predicted.
-We replace the traditional single SOS with the sequence-based SSOS.
-This approach incorporates the observed values from steps $0:\tau$ as inputs for both the encoder and decoder during the initial stage.
-During the autoregressive process, the outputs are progressively concatenated and fed back into the Trajectory Generator as inputs.
-This allows the model to predict the complete trajectory while computing the loss for all outputs (Equation 3).
+SSOS is designed to mitigate the divergence at the first step when using the Transformer’s Encoder-Decoder architecture for numerical time series prediction tasks (detail in Section 4.2), such as pedestrian trajectory prediction.
 To evaluate the effectiveness of SSOS, we introduce the Initial Displacement Error (IDE):
@@
 where $\tau+1$ denotes the first prediction time step.
-Table III shows the average ADE results across all datasets, comparing different ablations of our model.
-”w/o” refers to ”without” a component.
-The full model (InSyn) includes both RP and IS, while the Baseline model excludes both components.
-The results show that the SSOS strategy reduces the average IDE by approximately 6.58% compared to the SOS strategy.
+The results in Table 3 show that the SSOS strategy reduces the average IDE by approximately 6.58% compared to the SOS strategy.
 This reduction alleviates prediction divergence at time $\tau+1$, enabling smoother transitions between observed and predicted trajectories.
-Additionally, it mitigates error accumulation during autoregressive prediction.
-TABLE III: IDE Result of Ablation Study on SSOS
+Note that the improvement is more significant in Zara datasets, where pedestrians have fewer interactions and rely more on their own historical trajectories.
+The SSOS strategy, which reinforces the model’s encoding of individual motion patterns by integrating the observed sequence, thereby enhances self-expression and leads to greater gains.
+Table 3: IDE Result of Ablation Study on SSOS
 Methods IDE
@@
 InSyn-SOS 0.118 0.051 0.085 0.062 0.065 0.076
-V-C Case Study
+5.3 Case Study
 Figure 5: Case Study of Region Partition and Interaction State.
-This figure compares three variants of InSyn: (1) without region partition (w/o-RP), (2) without interaction state (w/o-IS), and (3) the full model incorporating both components (InSyn).
+This figure compares three variants of InSyn: (1) without region partition (w/o-RP), (2) without interaction state (w/o-IS), and (3) the full model (InSyn).
 To isolate the effect of interaction modeling, we exclude the Goal Sampler in this evaluation.
 To further investigate the effectiveness of region partition and interaction state, we conduct a case study.
-The cases (see Figure 5) cover various interaction scenarios, including paired walking, group walking, and pedestrian conflict.
-In Sync Scenarios.
+The cases extracted from the datasets (see Figure 5) cover various interaction scenarios, including paired walking, group walking, and pedestrian conflict.
+5.3.1 In Sync Scenarios.
 Cases (a), (b), (e), and (f) represent instances where pedestrians walk in coordination with others nearby.
@@
 Case (f) presents a simple, slow-moving scenario where all models achieve comparable performance, showing that the base model alone suffices for simple prediction tasks.
-Conflict Scenarios.
+5.3.2 Conflict Scenarios.
 Cases (c) and (d) illustrate interactions where conflict plays a dominant role.
@@
 This interpretability not only builds trust in the model’s decisions but also provides actionable insights for refining interaction modeling strategies.
-VI Conclusion
-This paper presents InSyn, a trustworthy trajectory prediction model that extends the Transformer architecture to explicitly capture pedestrian complex interaction patterns.
-By introducing interpretable structure into otherwise black-box models, InSyn improves interpretability and robustness in socially complex environments.
-Experimental results validate the effectiveness of our approach in both dense and sparse scenarios, offering more reliable predictions.
+6 Conclusion
+This paper presents InSyn, a trustworthy trajectory prediction model that extends the Transformer architecture to explicitly capture and leverage pedestrian complex interaction patterns.
+By introducing interpretable structure into otherwise black-box models, InSyn improves both accuracy and interpretability.
+Experimental results validate the effectiveness of our approach in both dense and sparse scenarios, offering more accurate predictions in complex interaction scenarios.
 Besides, our proposed SSOS strategy enhances prediction accuracy by mitigating initial-step divergence in numerical sequence modeling.
@@
 Future work will focus on two key directions.
-First, we aim to design a more refined approach for partitioning interaction regions to better handle challenges posed by small-scale scenes.
+First, we aim to design a more refined approach for interaction identification.
 Second, we plan to apply the SSOS strategy to other types of numerical sequential data beyond trajectory prediction, exploring its potential in diverse domains.
-Acknowledgment
-This work was supported by the Natural Science Foundation of Shanghai (No. 23ZR1422800). The authors would like to thank the SusCom Lab and Dr. Taowen Wang at HKUST(GZ) for their valuable support.